### v1

# Ethics at the Scalpel’s Edge: Trustworthy Surgical Video Analysis in the Era of Foundation Models

Abstract
Recent advancements in machine learning (ML) and deep learning (DL), particularly through the introduction of foundational models (FMs), have significantly enhanced surgical scene understanding within minimally invasive surgery (MIS). This paper surveys the integration of state-of-the-art ML and DL technologies, including Convolutional Neural Networks (CNNs), Vision Transformers (ViTs), and foundational models like the Segment Anything Model (SAM), into surgical workflows. These technologies improve segmentation accuracy, instrument tracking, and phase recognition in surgical endoscopic video analysis. The paper explores the challenges these technologies face, such as data variability and computational demands, and discusses ethical considerations and integration hurdles in clinical settings. Highlighting the roles of FMs, we bridge the technological capabilities with clinical needs and outline future research directions to enhance the adaptability, efficiency, and ethical alignment of AI applications in surgery. Our findings suggest that substantial progress has been made; however, more focused efforts are required to achieve seamless integration of these technologies into clinical workflows, ensuring they complement surgical practice by enhancing precision, reducing risks, and optimizing patient outcomes.
I Introduction
Artificial intelligence (AI) has the potential to have a profound impact in the field of surgery. This is particularly relevant in the field of “vision”, where improving the understanding of complex surgical scenes through advanced imaging and analysis techniques can complement surgical actions [1]. Minimally invasive surgery (MIS) has become the benchmark for advanced surgical procedures, offering numerous benefits over traditional open surgery, such as reduced trauma and quicker recovery times. These procedures rely heavily on the surgeon’s ability to interpret high definition (HD) video feeds that capture the dynamic and often unpredictable environment within the surgical field of view (SFOV). The complexity of these SFOV is characterized by variable lighting conditions, motion blur from rapid instrument movements, and hampered SFOV caused by blood, intracorporeal surgical “smoke” or other fluids. Traditional image processing methods often struggle with this variability, which limits their effectiveness in real-time applications[2].
Figure 1: Key components in surgical scene understanding: we explore the use of Surgical AI for Tool Detection and Tracking in Section III, Workflow Recognition in Section IV and Surgical Training and Simulation in Section V.
Figure 2: Organization of the review paper on surgical scene understanding categorized into seven sections, detailing the evolution of deep learning models in surgical applications, advancements in surgical tool recognition, and comprehensive analysis of surgical workflows, training, and simulation. The paper provides a structured roadmap through various technical discussions, highlights relevant datasets, and concludes with insights into open issues and future research directions in the domain of surgical scene understanding.
AI advancements such as deep learning (DL) have opened new possibilities for understanding surgical scenes. Unlike traditional methods that rely on predefined features, these AI models learn to recognize patterns and features directly from data, improving their ability to generalize across different settings. These capabilities are crucial for tasks such as distinguishing anatomical structures from surgical tools, navigate obscured views, and predict the presence of abnormalities or complications in real time [3]. In this context, the precision and accuracy of imaging and visualization technologies become paramount.
The development of state-of-the-art imaging techniques has significantly improved the ability of MIS surgeons to plan and execute complex surgeries with greater confidence and improved outcomes [4]. A crucial component of these advances is medical image segmentation, which involves partitioning digital images into distinct regions that represent different tissues, structures, or devices. Segmentation is essential for numerous medical applications, allowing improved visualization, perception, and interpretation of complex anatomical and pathological information [5, 6, 7]. In surgical contexts, effective segmentation ensures that surgical professionals accurately identify and differentiate between various anatomical features and surgical tools, thereby prompting the surgeon to precisely dissect in correct surgical planes thus avoiding collateral damage to adjacent structures. For example, the left ureter can sometimes be damaged when releasing the sigmoid colon off the retroperitoneal structures in rectal cancer surgery (anterior resection). This focus on developing surgical safety systems, analogous to “satellite navigation”, has the potential to highlight potential hazards, minimize complications and therefore improve patient outcomes.
The key tasks of surgical scene understanding as shown in Fig. 1, are outlined next.
  • Surgical Tool and Object Detection and Tracking: Identifying anatomical objects and following the movement of instruments within the surgical field provides context-aware assistance and ensures patient safety. For example, identifying the proximity of a surgical instrument close to a blood vessel or adjacent organ would enhance situational awareness within the surgical field [8].
  • Surgical Workflow Recognition: Identifying different phases of a procedure enables real-time documentation, skill assessment, and context-aware assistance [9]. Additionally, recognizing surgeon gestures and actions supports training, performance evaluation, and automation. This serves as a foundation for converting surgical videos into structured formats, enabling large-scale indexing and advanced data analysis [10].
  • Surgical Training & Simulation: Understanding the surgeon’s gestures and actions can be used for training, assessment, and automation purposes [10]. Detecting unexpected events (for example, bleeding or damage to adjacent tissue or organs) [11], generating synthetic videos [12], and performing question-answering feedback tasks [13] can also lead to enhanced surgical experience from the perspective of clinicians.
Figure 3: An in-depth examination of the intricate tasks involved in navigating and controlling surgical tools, highlighting the technical and operational processes that enable precision and effectiveness in modern surgical procedures
I-A Comparison with Related Surveys
DL applications in medical imaging and surgery have been extensively reviewed, with numerous studies summarizing advancements and identifying emerging trends. This survey provides a different perspective by focusing on the predominant areas and methodologies shaping the domain of understanding surgical video scenes. In Table I, we present a comprehensive comparison of this article with existing surveys and review articles that have a similar focus. Specifically, a comparison is provided in terms of scope, procedural coverage, datasets utilized, methodological approaches, algorithmic details, performance metrics, and unique contributions.
Compared to the existing literature, this survey adopts a more comprehensive and integrative approach and attempts to address various methodologies and applications within the domain of understanding the surgical scene. Previous works, such as those of Rivas et al. [14] and Rueckert et al. [15], have predominantly focused on specific tasks such as automation in MIS and instrument segmentation, respectively. In contrast, this survey expands the focus to encompass a broader spectrum of challenges, including segmentation, tracking, and workflow recognition. Moreover, unlike the studies by Fu et al. [16] and Azad et al. [17] that mainly focuses on endoscopic navigation and U-Net variants, respectively, and explore individual methodologies, this paper emphasizes the applicability of advanced models such as Vision Transformers (ViTs), Large Vision-Language Models (LVLMs), and foundation models like SAM. These models are particularly highlighted for their potential to assist and enable real-time decision making and address the complexities associated with multimodal data. Furthermore, based on the work of Li et al. [18], this survey enhances existing data set analyzes and presents an updated and detailed catalog specifically tailored for surgical scene understanding tasks. By synthesizing these contributions, this survey not only consolidates current knowledge but also identifies key trends, such as the integration of multimodal AI and foundation models for surgical applications. It also underscores the pressing need for real-time decision-making tools that meet the dynamic and “high-stake” demands of live surgical environments, providing a robust framework to advance AI-driven innovations in surgical practice.
TABLE I: Comparative performance and limitations of DL-based medical image segmentation surveys.
Reference Scope Contributions Algorithms Generative Models Vision-Language Models Foundational Models Ethical and Regulatory Insights Dataset Coverage Comprehensive Surgical Tasks Foundational Models Adaptation Clinical Perspective Open Issues
Rivas-Blanco et al. [14] DL models in minimally invasive surgery Deep insights into automation of surgical tasks CNN, RNN, ANN, HMM
Fu et al. [16] Endoscopic navigation technologies Detailed exploration of endoscopic vision technologies Augmented reality, various imaging modalities
Rueckert et al. [15] Instrument segmentation in surgery Comprehensive review on surgical instrument segmentation Deep learning, CNN
Zhang et al. [19] SAM in medical image segmentation Exploration of SAM’s extension to medical segmentation SAM, foundation models
Azad et al. [17] Evolution and success of U-Net in medical imaging In-depth analysis of U-Net variants across modalities U-Net and its variants
Schmidt et al. [20] Tracking and mapping in medical CV Insightful review of tracking/mapping in deformable tissues Nonrigid tracking, SLAM
Upadhyay et al. [21] DL methods to overcome data scarcity Extensive review on overcoming data scarcity CNN, U-Net, GAN
Zhou et al. [22] AI in surgery integration Comprehensive integration of AI in surgery CNN, RNN
Morris et al. [23] DL applications in sub-specialties Insights into practical DL applications in surgery CNN, RNN
Li et al. [18] Surgical workflow recognition Detailed surgical workflow recognition analysis CNN, LSTM
Garrow et al. [24] Automated phase recognition Highlight on phase recognition automation in surgery HMM, ANN
This Paper Surgical Scene Understanding Detailed model review, applications overview, dataset analysis CNNs, ViTs, LVLMs, GANs, FMs
I-B Contributions of This Paper
This survey provides a focused and contemporary review of understanding of the surgical scene based on AI, emphasizing the transformative role of foundation AI models in the field. Unlike traditional reviews that primarily cover conventional ML and DL techniques, we delve into the unique advancements enabled by foundation models, Vision Transformers (ViTs), pre-trained multimodal architectures, and generative AI systems. Our contributions are as follows:
  • Contemporary Focus on Foundation AI Models: This survey uniquely highlights how foundation models are reshaping surgical scene understanding, particularly through their ability to generalize across tasks and modalities. We analyze their applications in real-time surgical workflows for tasks such as robust tool segmentation, fine-grained workflow recognition, and anomaly detection, emphasizing their scalability and adaptability in MIS.
  • Critical Insights into Transformative Use Cases: By focusing on endoscopic video analysis, we underscore how foundation models outperform traditional approaches in addressing challenges like variability in surgical scenes, occlusions, and inter-patient heterogeneity. The survey provides critical insights into their utility in tasks such as automated annotation, video summarization, and cross-modal reasoning, bridging gaps in surgical data analysis with unprecedented efficiency.
  • Evaluation of Evolving Datasets and Benchmarks: We offer an in-depth evaluation of emerging datasets and benchmarks specifically designed for foundation model training and validation in surgical applications. This includes an analysis of large-scale, multimodal datasets that enable transfer learning and fine-tuning for surgical scene understanding, addressing gaps in data availability and variability.
By centering on foundation AI models, this survey represents a timely and significant advancement in the literature, offering a forward-looking perspective on their role in revolutionizing surgical practices. It serves as a critical resource for researchers and clinicians aiming to leverage state-of-the-art AI to enhance surgical scene understanding in MIS, a fundamental step in improving patient outcomes.
I-C Organization of This Paper
This paper provides a comprehensive review of foundational AI models and their application in surgical scene understanding, with a particular focus on interpreting endoscopic video data for minimally invasive surgery. By exploring the integration of advanced AI techniques, this review highlights the transformative impact of these technologies on surgical practice, with the aim of improving precision and outcomes in MIS. The review also delves into the evolution and current landscape of various techniques within medical imaging, including segmentation, tool detection, workflow recognition, and anomaly detection, specifically their utility in surgical environments. It traces the chronological development of these technologies and examines contemporary advancements, emphasizing their critical role in augmenting surgical precision and improving patient outcomes. The organization of the paper is shown in Fig. 2, which provides the visual presentation of all the sections and sub-sections.
II Background and Challenges
II-A Surgical Video Analytics and Scene Understanding
Surgical video analytics and scene understanding represent transformative advances at the intersection of computer vision, ML, and surgical practice, with the aim of improving the safety, efficiency, and precision of surgical procedures. These technologies leverage cutting-edge techniques from image processing and ML / DL to analyze videos captured during surgery, providing actionable insights in real time and impact on postoperative events (for example, increased bleeding events may lead to increased length of hospital stay; increased intraoperative bowel handling may lead to bowel paralysis and higher rates of post-operative nausea and vomiting). Algorithms for detecting, segmenting, and tracking surgical instruments and maneuvers transform complex video data into precise information that supports decision making, as illustrated in Fig. 3.
Surgical scene understanding complements video analytics by focusing on the interpretation of surgical environments, including instrument recognition, workflow analysis, and scene segmentation. These capabilities enable real-time decision support through augmented overlays that guide procedural steps, comprehensive postoperative analyzes to assess surgical efficacy, and automated documentation to reduce administrative burden on surgical teams, particularly relevant after a procedure that may take several hours. The latter would allow clinicians to focus on other clinical tasks such as informing relatives of patients about the procedure or assisting in the safe transfer of patients out of the operating room. In addition, it is likely to describe the ground truth of operative events without bias, a highly relevant point in governance. In Fig. 4, the three stages of surgical procedures are depicted, illustrating the comprehensive workflow from pre-operative planning through intra-operative execution to post-operative recovery and evaluation.
Furthermore, the advent of 3D augmentation technologies has revolutionized surgical video analytics by converting traditional 2D video streams into dynamic 3D reconstructions, providing surgeons with an immersive and intuitive view of the surgical site [25]. These systems improve spatial awareness, simplify navigation around critical anatomical structures, and enable more precise interventions [26], minimizing patient trauma and improving outcomes.
Despite these advances, numerous challenges persist, including variability in surgical procedures, differences in patient anatomy, and the unstructured nature of surgical environments, which collectively complicate algorithm development. Moreover, high-dimensional video data demands substantial computational resources for real-time processing, while stringent privacy and security measures remain critical to safeguarding sensitive medical footage. The future of surgical video analytics and scene understanding lies in deeper integration with AI and multimodal data, combining video analysis with patient vitals and other intraoperative information. Innovations like edge computing promise real-time analytics in resource-constrained settings, whereas enhanced algorithms can provide greater granularity and accuracy in data interpretation. These advancements are poised to redefine surgical precision, training, and patient care, ultimately paving the way for safer and more effective surgical practices worldwide.
Figure 4: Sequential stages of a surgical procedure. The Pre-Operative stage involves patient assessment, detailed imaging (and scene understanding), and surgical planning (based on better understanding). The Intra-Operative stage focuses on the actual surgical procedure itself, including anesthesia administration, correct patient positioning and adherence to safety protocols. The Post-Operative stage covers both initial and long-term recovery, concluding with follow-up and outcome evaluations to ensure the effectiveness of the procedure.
II-B DL Models in Surgical Video Analysis
II-B1 Convolutional Neural Networks
CNNs are foundational to modern video analysis, particularly due to their ability to efficiently extract high-level features from visual data. Structured as a series of convolutional layers, CNNs capture spatial hierarchies in images, rendering them extremely effective in tasks such as object detection and scene classification. These networks perform convolutions across image pixels to generate feature maps that summarize the presence of specific features at various locations in the image. This ability makes CNNs particularly adept at processing visual inputs [27] that are common in video data, such as frames from surgical procedures. In surgical video analysis, CNNs are utilized not only for object identification but also for understanding the interaction and relative positioning of various surgical tools and anatomical structures. For example, through frame-by-frame analysis of video data, CNNs can detect environmental changes, track the movement of surgical instruments, and monitor the progress of surgical interventions and any potential hazards, and predict the end of the surgical procedure (this helps in high-level management of theater lists and helps improve case throughput and theatre time utilization). Such capabilities are crucial for developing real-time feedback systems that can alert surgeons about critical events or anomalies detected during surgery, as well as giving theatre teams a high-level overview of surgical procedures [28].
II-B2 Segmentation Models
Segmentation models are crucial in medical image analysis, providing detailed pixel-level annotations that are essential for both diagnostic and interventional procedures. Two of the most influential segmentation models in this domain are U-Net and DeepLabv3, each known for its effectiveness in handling the complexities of medical imagery.
  • U-Net: Developed specifically for biomedical image segmentation, U-Net features a symmetric architecture with a downsampling path to capture context and a precisely corresponding upsampling path to allow for precise localization [29]. This network architecture is particularly well suited for medical applications because of its efficiency in using a limited number of training samples to produce high-resolution segmentations. The ability of U-Net to perform well even with small amounts of data and its robustness against noise in images make it a preferred choice for segmenting surgical video frames, where precise delineation of organs and surgical instruments is critical.
  • DeepLabv3: Building on earlier versions, DeepLabv3 incorporates atrous convolutions, a technique that increases the receptive field of filters by inserting spaces between kernel elements, to capture richer contextual information without compromising the sharpness of the image segmentation [30]. This model utilizes an atrous spatial pyramid pooling module to robustly segment objects at multiple scales, a feature particularly useful for the varied scales seen in surgical videos. DeepLabv3’s ability to effectively segment fine structures at different depths and scales is invaluable in surgeries, providing clear delineations that can guide surgical interventions and improve outcomes.
Figure 5: Overview of foundational models applied in surgical settings, illustrating their roles in various applications such as report generation, treatment planning, and surgical assistance.
II-B3 Vision Transformers (ViT)
ViT represents a significant paradigm shift in how image data is processed for complex tasks like video analysis, including in high-stakes environments such as operation theaters. Originally adapted from the transformer architecture, which has revolutionized natural language processing, ViTs apply the principles of self-attention to visual contexts, allowing them to learn contextual relationships between different parts of an image [31]. In surgical video analysis, ViTs have shown great promise in accurately segmenting and identifying critical structures within surgical scenes, outperforming traditional CNNs in tasks that require understanding complex spatial dependencies and long-range interactions [32]. This makes ViTs particularly suitable for applications where precision and context-aware decision-making are crucial, such as real-time surgical guidance and post-operative analysis.
II-B4 Diffusion Models
Diffusion models are a class of generative models that have attracted significant attention for their ability to generate high-quality, detailed images from a learned distribution of training data. These models work by initially learning to gradually add noise to an image, transforming it into a Gaussian noise distribution, and then learning to reverse this process to reconstruct the original image from the noise [33]. This forward and reverse process enables diffusion models to effectively model the probability distribution of training data, making them extremely powerful for generating or reconstructing images. In surgical video analysis, diffusion models can be utilized to enhance low-resolution or noisy surgical videos, providing clearer visualizations of surgical procedures . They can also generate synthetic surgical images for training purposes, allowing for the creation of diverse scenarios that help train surgical staff without the need for extensive real-life video collections. Additionally, diffusion models, with their generative capabilities, hold significant potential for simulating possible surgical outcomes based on intra-operative videos, aiding surgeons in planning and decision-making during complex procedures. These applications demonstrate the versatility and potential of diffusion models in transforming surgical education for future generations of surgical trainees and practice by enhancing the quality and utility of surgical imagery. Ultimately, a library of surgical videos that create all potential scenarios or hazards that can occur during surgery will lead to better prepared surgeons and improved patient outcomes.
II-C Foundation Models
Foundation models (FMs) represent a significant breakthrough in machine learning, characterized by their extensive pre-training on large-scale datasets across a diverse range of tasks and domains [34]. Although these models are generally designed for broad use cases, researchers are increasingly applying and fine-tuning them in specialized fields like medical imaging. In the realm of medical imaging, FMs have revolutionized the approach to complex image data interpretation. These models leverage the massive amount of pre-training to develop a deep and nuanced understanding of image features, which is crucial for tasks such as medical image segmentation, classification, and anomaly detection. The application of these models in medical image analysis has led to substantial advancements, particularly in enhancing the accuracy and efficiency of segmentation processes used to identify and delineate anatomical structures and pathological conditions from medical imaging data [35].
Despite these benefits, there are notable challenges when using FMs in healthcare. First, patient privacy must be upheld, which can limit the volume and variety of data available for training [36]. Second, the “black box” nature of deep learning can make it difficult to interpret model decisions—an issue of particular concern in medical settings, where transparency and accountability are paramount [37]. Third, many clinical environments lack the substantial computational resources or extensive datasets needed to fine-tune large models [38]. Further complicating matters are rigorous data protection laws, such as HIPAA in the United States and GDPR in the European Union, which limit data sharing across organizations and regions.
Nevertheless, FMs also open exciting new possibilities. They can integrate multiple data types, such as combining medical images with patient health records to create more comprehensive diagnostics [39]. By refining these models through detailed customization, it is possible to address specific needs in areas like surgical video analysis and disease detection. They introduce a paradigm shift in medical image analysis by leveraging large-scale datasets to capture a broad spectrum of features and patterns [40]. Ultimately, while FMs hold great promise, ongoing research is needed to adapt them effectively to diverse medical tasks, ensure interpretability, and comply with strict regulatory standards [41].
Finally, while foundation models hold great promise for broad applications, their effectiveness in specialized medical contexts such as surgical video analysis and diagnostics relies heavily on customization and domain-specific fine-tuning [42]. This section examines models like SAM, CLIP, DINO, BERT, and GPT, which showed instrumental outcomes in advancing data processing and decision-making in the medical domains, as shown in Fig. 5.
II-C1 Segment Anything Model (SAM)
The SAM [43] was developed to perform general-purpose image segmentation with minimal task-specific retraining using over 11 million images. Its adaptability extends to a wide array of applications, including autonomous driving, medical imaging, and satellite image analysis. Despite its non-medical origin, SAM shows considerable promise in healthcare settings. By combining robust CNN features and deep learning techniques, SAM can quickly adapt to different segmentation tasks, such as identifying tumors in CT scans, surgical tool recognition, or delineating organs in MRI data. This ability is referred to as zero-shot segmentation, as it allows SAM to perform inference on new imaging problems without extensive domain-specific training.
The subsequent iterations and specialized adaptations of SAM further illustrate its effectiveness and flexibility. For example, SAM3D adapts its architecture for 3D imaging tasks (e.g., CT, MRI, PET), using specialized 3D CNN layers [44]. In SAM, further innovations have been introduced that integrate multimodal imaging data, such as MRI and ultrasound, to achieve comprehensive diagnostic precision [45]. Further adaptations have been developed for the medical domain, such as the Medical SAM Adapter (MSA), which fine-tunes the SAM using adapter layers for higher segmentation precision in medical contexts [46]. SAM2 [47], which is a recent advanced version of SAM, has shown significant improvements in performing real-time segmentation in videos, which demonstrates its enhanced capabilities in complex scenarios such as tumor identification and organ delineation [48]. Lastly, many different efforts are being made to incorporate domain-specific knowledge into these models to capture the nuances of medical images more effectively using limited data [49].
(a) Instrument flare
(b) Partial occlusions by blood
(c) Smoke-induced occlusion
(d) Underexposed regions
(e) Motion blur
(f) Multiple instruments
(g) Partial occlusion by organ
(h) Transparent instrument
Figure 6: Key challenges in surgical video analysis include photometric artifacts (e.g., blurriness, specular reflections), occlusions from blood and tissue deformation, limited camera view, and tool similarity. These complexities hinder robust localization and segmentation, emphasizing the need for advanced foundation models. Image adapted from [50].
II-C2 Generative Pre-trained Transformer (GPT)
The GPT [51] foundation model, developed by OpenAI, serves as a versatile backbone for numerous AI applications due to its exceptional generalization and natural language understanding capabilities. ChatGPT is the most prominent development by OpenAI, which is an instruction-tuned conversational model that adapts GPT’s capabilities for interactive and user-centric applications. While GPT functions as a general-purpose foundation model capable of handling a wide range of language tasks, ChatGPT is fine-tuned to specialize in conversational contexts, making it particularly effective in medical settings. It enables the automation of routine inquiries, enhances patient engagement through interactive dialogues, and streamlines clinical workflows by converting unstructured data into accurate and detailed medical reports. Moreover, its ability to assimilate and articulate complex medical literature makes it an invaluable tool for supporting research and continuing education for healthcare professionals. ChatGPT’s integration into clinical practice not only improves operational efficiency but also facilitates personalized patient care by providing timely and relevant medical information. However, deploying ChatGPT in healthcare requires rigorous attention to data privacy, ethical considerations, and model reliability to ensure secure and responsible use in sensitive medical environments [52].
II-C3 Bidirectional Encoder Representations from Transformers (BERT)
BERT is a pioneering language model that has significantly advanced natural language processing through its ability to understand context bidirectionally using a transformer-based architecture [53]. In the surgical domain, the complexity and specificity of medical language require sophisticated models such as BERT, which are adept at capturing the nuanced semantics and intricate terminologies inherent in surgical texts. BERT’s pre-training on extensive and diverse datasets enables it to develop a deep comprehension of language patterns, making it well-suited to address the rigorous demands of surgical data analysis [54]. Its architecture facilitates an effective contextual representation, which is critical for accurately interpreting the multifaceted information present in surgical documentation and communication. Moreover, the adaptability of BERT through fine-tuning processes ensures that it can be tailored to meet the specific linguistic and contextual requirements of the surgical field, thereby enhancing the robustness and precision of language understanding in medical contexts. Although primarily focused on textual data, BERT also supports the integration and interpretation of findings from various medical imaging modalities, such as MRI or CT, by aiding in the automatic extraction and structuring of textual reports generated from these images, thus bridging the gap between radiological findings and surgical planning. This would be hugely useful in large-scale rapid automated reporting and would help alleviate the pressures human radiologists face when reporting large volumes of scans under the Faster Diagnostic Standards (FDS). In the UK, the FDS framework is a 28-day national health service (NHS) target to diagnose or exclude cancer in patients. These attributes position BERT as a foundational model in integrating advanced language processing capabilities within surgical data systems, contributing to the advancement of data-driven practices in surgical environments.
II-C4 Self-Distillation with No Labels (DINO)
DINO [55] exemplifies a groundbreaking approach in self-supervised learning, initially designed for general computer vision tasks. Nevertheless, its ability to learn from unlabeled data makes it especially appealing for medical applications, where annotated data can be hard to obtain. By employing a teacher-student training framework, DINO captures meaningful image features from different augmented views of the same input without needing pre-existing annotations. This is very important for the medical domain as the data is often scarce and hard to annotate.
In surgical contexts, for example, DINO could learn to recognize and segment important structures in endoscopic videos even if ground truth labels are scarce. This feature is crucial in clinical environments, where the time and expertise required to annotate large volumes of surgical footage may be limited. Many researchers have utilized DINO for different applications in several innovative ways. For instance, the SurgVID [56] framework uses DINO’s self-supervised learning to efficiently segment surgical tools and anatomical structures from video data, thus reducing the need for extensive manual annotations. This approach has proven to be nearly as effective as fully-supervised methods, showcasing the potential of self-supervised techniques in surgical applications.
Another adaptation, known as SurgicalDINO [57], modifies DINO for depth estimation in robotic surgery. This version incorporates Low-Rank Adaptation (LoRA) layers [58], allowing the model to be fine-tuned for specific surgical applications without the need to retrain the entire model. This method demonstrates how DINO can be tailored to surgical navigation and 3D reconstruction. Additionally, Ramesh et al. [59] compared various self-supervised learning models, including DINO, on surgical datasets like Cholec80. They explored how well these models perform in recognizing surgical phases and detecting tools, indicating that self-supervised learning holds promise for enhancing surgical computer vision.
Furthermore, DINO’s scalability and efficiency make it well-suited for medical centers that lack extensive computational resources. Through these advantages, DINO holds considerable promise in supporting tasks like surgical planning, intraoperative guidance, and postoperative analysis.
II-C5 Large Language Model Architecture (LLaMA)
LLaMA [60], a high-capacity language model, leverages transformer-based mechanisms to understand and generate human language with remarkable accuracy and depth. Developed to facilitate advanced natural language processing tasks, LLaMA is designed for a wide array of applications ranging from automated text generation to complex query handling [61]. In the medical field, LLaMA’s capabilities are particularly beneficial, as the model can interpret and synthesize medical literature, patient reports, and clinical guidelines with high precision. By training in various medical texts, LLaMA can help healthcare professionals by providing diagnostic suggestions, summarizing patient histories, and even generating informational content for patient education. Its ability to process and generate medical text effectively makes it an invaluable tool for enhancing clinical decision-making and improving patient outcomes. Moreover, LLaMA integration into clinical information systems can streamline workflows by automating documentation processes and extracting useful information from vast datasets, allowing medical personnel to focus more on patient care than administrative tasks [62]. As the demand for efficient and accurate processing of medical information grows, LLaMA stands out as a transformative model capable of revolutionizing various aspects of healthcare delivery.
II-C6 Contrastive Language-Image Pre-training (CLIP)
CLIP [63] is a novel model introduced by OpenAI that revolutionizes the way machines understand images and text together. CLIP is trained in a contrastive manner where the images are paired with the corresponding text, learning from a wide variety of publicly available images and captions. This unique training approach allows CLIP to generalize across a wide array of visual concepts in a zero-shot manner. It can understand and categorize images that it has never seen before based on textual descriptions alone [64]. In the medical and surgical fields, CLIP’s capabilities can be particularly transformative. Its ability to interpret and correlate complex medical imagery with corresponding clinical notes or annotations without direct supervision makes it an excellent tool for diagnostic imaging. For example, CLIP can help radiologists and surgeons quickly identify relevant features in medical scans, such as magnetic resonance images or CT images, that correspond to textual descriptions found in case reports or diagnostic criteria. This could significantly speed up the diagnosis process, an imperative in the NHS FDS pathway, and improve the accuracy of identifying and classifying pathological characteristics. Moreover, CLIP’s robust generalization ability allows it to adapt to diverse medical datasets, potentially reducing the time and resources required for model training and fine-tuning in specialized medical applications[65].
II-D Challenges in Surgical Scene Understanding
Surgical scene understanding presents a diverse array of challenges that stem from the complexity and unpredictability of surgical environments, coupled with the technical limitations of current AI models. To advance AI-assisted surgical systems, the following challenges must be carefully addressed.
II-D1 Technical Challenges
The domain of surgical video analysis presents unique challenges that hinder robust localization and segmentation of surgical instruments, as shown in Fig. 6. Unlike natural images and videos, surgical frames are characterized by high tissue deformations and frequent occlusions caused by the presence of blood and multiple artifacts on the instruments. Photometric artifacts, as identified by [66], can significantly degrade the performance of segmentation models. Some of the additional complexities include:
  • a)
    Subtle Variance and Limited View: Surgical procedures may involve subtle interphase or intraphase variances that are difficult to capture consistently. The limited field of view offered by surgical visualization cameras further complicates visual assessment, while restricting the visual context available for decision-making [67].
  • b)
    Blurriness and Specular Reflection: Camera motion and the gas emissions from surgical tools often cause blurriness, while specular reflections and scale variations, as discussed by Baumhauer et al. [68], can lead to poor segmentation accuracy.
  • c)
    Tool Similarity and Edge Presence: The segmentation of multiple instruments is particularly challenging due to the appearance and shape similarity between different tools. Instruments located on the edge of video frames are especially difficult to detect and segment reliably [69]. Moreover, the variations in the instrument pose may also alter the perceived geometry or shape, depending on the surgical camera’s field of view.
TABLE II: Overview of Machine Learning Applications in Surgery
Application Method Model Data Metrics Limitations
Instrument Segmentation [29] DL U-Net Surgical videos Accuracy, IoU High computational cost
Tissue Classification [70] Supervised Learning CNNs Histopathological images Precision, Recall Data variability
Phase Recognition [71] Temporal Analysis LSTM Time-series data Accuracy, F1-Score Long training times
Anomaly Detection [72] Anomaly Detection Autoencoders Physiological signals AUC, ROC False positives
Skill Assessment [73] Pattern Recognition SVM Motion tracking data Accuracy, Error rate Subjective bias
Organ Segmentation [74] DL V-Net CT scans, MRI Dice coefficient Requires large datasets
Error Prediction [75] Predictive Modeling Random Forest Operational data Sensitivity, Specificity Overfitting
Depth Estimation [76] Computer Vision Stereo Vision 3D imaging Mean error Calibration issues
Segmentation [15] DL CNNs Endoscopic Videos Accuracy, IoU Computational demand
Medical Image Segmentation [19] Deep Learning, Foundation Models SAM CT, MRI Dice Coefficient, IoU Poor zero-shot performance, requires fine-tuning
Medical Image Segmentation [17] DL U-Net CT, MRI, US, X-ray, OCT, PET IoU High computational cost, limited data
Biomedical Image Segmentation [77] Deep Learning, Zero-Shot Learning SAM Medical images across various modalities Accuracy, Dice Coefficient, IoU Segmentation in complex anatomical regions, adaptation to domain-specific challenges
Tracking and Mapping [20] Computer Vision, SLAM Various, including SLAM and SfM Medical Imaging (Endoscopy, Colonoscopy) Accuracy, Precision Low texture, light reflections, organ deformation
3D Reconstruction and Localization [78] Computer Vision Laparoscope Localization Surgical Videos Accuracy, Precision Needs high computational power, complex setup
3D vs. 2D Vision in Laparoscopy [79] Systematic Review Not Applicable Surgical Training Performance Time, Error Reduction High cost, visual discomfort, fatigue
Surgical Tool Detection [8] Review, Machine Learning Not Applicable Surgical Tools Accuracy, Recall, Precision Requires robust feature extraction and tracking methods, sensitive to visual occlusions and tool coverage
Endoscopic Video Processing [80] Review Various, Computer Vision Endoscopic Videos Accuracy, Precision Requires robust computational techniques, sensitive to video quality and consistency
Tracking [8] Comparative Study, DL CNNs, RF Surgical videos Accuracy, Precision, Recall Complexity of dynamic environments, inconsistent annotations
II-D2 Lack of Representative Datasets
The application of foundational models in surgical scene understanding faces significant challenges due to the scarcity of labeled data essential for training supervised learning methods. This scarcity is exacerbated by class imbalance issues, notably between foreground (surgical instruments) and background, where instruments occupy fewer image pixels compared to the predominantly background pixels. Current state-of-the-art approaches focus on pixel-wise classification for instrument segmentation but fail to account for global semantic correlations across images, potentially leading to inaccurate feature distributions and compromised segmentation accuracy. To overcome these challenges, it is imperative not only to enhance the robustness and efficiency of AI models, but also to innovate in data set creation and enhancement. This will ensure the availability of high-quality and balanced data that is necessary to effectively train foundational models in the complex setting of surgical operations.
II-D3 Ethical, Regulatory, and Privacy Challenges
In addition to technical barriers, significant ethical, regulatory and privacy concerns must be addressed when applying AI to the understanding of the surgical scene. These considerations are essential to ensure that AI systems remain safe, reliable, and equitable in clinical settings. Below are some of the key challenges that require attention:
  • a)
    Patient Privacy and Data Protection: The use of AI in surgical procedures frequently involves sensitive patient data, raising major concerns about privacy and data security. The protection of this information is vital to maintain trust among patients, clinicians, and institutions [81].
  • b)
    Regulatory Compliance and Bias: AI systems in healthcare must comply with stringent regulatory standards, which continue to evolve [82]. Furthermore, AI models may introduce bias, especially if they are trained on nonrepresentative datasets, leading to disparities in care across different patient groups.
  • c)
    Surgeon Accountability and Trust: The introduction of AI into surgical decision making requires clear guidelines on surgeon accountability [83]. Although AI can assist in real-time decision making, surgeons must retain ultimate responsibility for patient outcomes. Ensuring that AI systems act as supportive tools rather than autonomous decision makers is critical to preserving trust and ethical responsibility.
III ML/DL Applications in Surgical Tool Recognition
The integration of ML technologies into surgical practices has catalyzed significant advances in computer-assisted interventions, improving both the precision and efficiency of procedures. These technologies facilitate a variety of applications, from diagnostic imaging to real-time operational guidance, fundamentally transforming the surgical landscape. The systematic deployment of various ML methodologies has great potential in improving surgical outcomes as well as streamlining both preoperative planning and postoperative care. As detailed in Table II, ML applications in surgery use a diverse range of techniques, including supervised learning for tissue classification, unsupervised anomaly detection, and DL for intricate tasks such as instrument segmentation and surgical skill evaluation. These techniques employ various models such as CNN for image-based analysis, SAM for zero-shot segmentation, and RNN for time series data, addressing specific needs within the surgical workflow. Table II categorizes these applications by specifying the ML methods used with the corresponding models and the data types utilized, ranging from histopathological images to surgical video data. It also outlines the metrics used to evaluate performance, such as accuracy, precision, recall, and the intersection over union (IoU). In particular, the table acknowledges the limitations inherent to each application, such as the high computational costs and the need for extensive data annotation, which are critical considerations for future research directions.
TABLE III: Comprehensive Summary of Tool Navigation Datasets (Chronologically Sorted)
Year Collection Data Volume Surgical Type Access Instrument Type Label Types Operational Tasks
2015 FetalFlexTool 21 Videos Fetal Surgery Public Rigid Bounding-box Detection
2015 NeuroSurgicalTools 2476 Images Neurosurgery Public Robotic Bounding-box Detection
2015 EndoVis 15 9K Images Colorectal Surgery Public Rigid Pixel-wise, 2D pose Segmentation, Tracking
2016 Cholec80 80 Videos Cholecystectomy Public Rigid Bounding Box Detection, Activity, Skills
2016 M2CAI16-tool 16 Videos Cholecystectomy Public Rigid Tool presence Detection, Phase
2017 Hamlyn 2 Phantom Cardiac Public NA Depth map Tracking
2017 ATLAS Dione 8 Videos In-vitro Experiments Public Robotic Bounding Box Detection, Activity
2017 EndoVis 17b 10 Videos Porcine Public Robotic Pixel-wise Segmentation, Binary, Parts
2018 m2cai16-tool locations 16 Videos Cholecystectomy Public Rigid Bounding-box Detection
2018 LapGyn4 55K Images Gynecologic Surgery Public Rigid No annotation Multiple
2018 EndoVis 18d 14 Videos Nephrectomy Public Robotic Pixel-wise mask Scene Segmentation
2019 SCAREDf 27 Videos Porcine NA NA Depth + mask Depth estimation
2019 Cata7 7 Videos Cataract Surgery Private Rigid Pixel-wise mask Segmentation, 3D Reconstruction
2019 UCL 16016 Synthetic Images Sigmoid resection, colonoscopy NA NA Depth map Depth estimation
2019 ROBUST-MIS19 30 Videos Proctocolectomy, rectal Public Rigid Instances Segmentation(Binary,Parts)
2020 LapSig300 300 Videos Colorectal Surgery Private Rigid Pixel-wise mask Segmentation, Parts, Action Recognition
2020 Sinus Surgery-L 3 Videos Sinus-Live Public Rigid Pixel-wise mask Segmentation, Phase, Action Recognition
2020 Sinus Surgery-C 10 Videos Sinus-Cadaver Public Rigid Pixel-wise mask Segmentation(Binary)
2020 UCL dVRK 20 Videos + Kinematic Data Ex-Vivo Public Robotic Camera parameters Segmentation(Binary)
2020 HeiSurF 24 Videos, 9 Test Surgeries Laparoscopic gallbladder resections Public Rigid Segmentation masks, Instrument classes Full scene segmentation, Workflow analysis
2021 KvasirInstrument 590 endoscopic frames Gastrointestinal Endoscopy Public Rigid Binary masks, Bounding boxes Tool Segmentation
2021 RoboTool 514 Video Frames, 14720 Synthetic Images Various Surgical Public Robotic Binary labels, segmentation masks Synthetic dataset creation
2021 I2I Translation 200,000 Images Laparoscopic Surgery NA Rigid Style transformation masks Image style translation
2021 SCARED 9 Datasets Porcine Public Robotic Depth Depth estimation
2021 CaDTD 50 Videos Cataract Surgery Private Rigid Bounding-box Detection
2021 dVPN 48702 Images Nephrectomy NA NA TP, A, SC, Ph Detection, Action Recognition
2021 EndoVis 21 33 Videos Cholecystectomy Public Rigid TP, A, SC, Ph Detection, Phase, Action Recognition
2021 AutoLaparo 21 Videos Laparoscopic Hysterectomies Public Rigid Annotated for uterus and instruments Workflow recognition, Motion prediction, Anatomy segmentation
2021 ART-Net 29 Procedures Laparoscopic Hysterectomies Public Non-robotic Binary segmentation, tool presence Instrument segmentation, geometry annotation
III-A Instrument Segmentation
Instrument segmentation is crucial in medical image analysis, especially in robotic-assisted surgery, where it involves accurately differentiating surgical instruments from the background and other elements within the scene. The effectiveness of robotic surgery systems hinges on precise visual information provided by segmentation algorithms, primarily based on CNNs. These algorithms process images in real-time to deliver clear outlines of surgical instruments, which is essential for tasks like automated tool tracking and interaction handling in robotic surgeries.
A significant contribution to this field is detailed by García-Peraza-Herrera et al. [84], who explore CNN-based methods for real-time surgical tool segmentation in laparoscopic videos. This study underscores the improvements DL models bring to surgical tool segmentation, improving clarity and operational efficiency in robotic surgeries, which are critical to ensuring safety and precision.
In Table IV, the advanced models are listed for instrument segmentation, covering a spectrum from supervised learning with models like FCN and U-Net to complex methods involving adversarial networks and domain adaptation. These models are evaluated on datasets like EndoVis and various private datasets, utilizing techniques that combine CNNs with RNNs and residual networks.
Transitioning from traditional segmentation techniques, the Min-max Similarity model introduces a novel contrastive and semi-supervised learning framework that marks a significant advancement in medical imaging. This model addresses the challenge of limited annotated data [85] by maximizing the similarity between representations of similar objects while minimizing those of dissimilar ones. The innovative approach improves the ability to distinguish between surgical tools and complex backgrounds, leveraging unlabeled data effectively to overcome the scarcity of labeled datasets. Its semi-supervised nature facilitates generalization across different surgical tools and procedures, reducing the need for extensive retraining. This method not only advances technical capabilities, but also provides substantial clinical benefits by enhancing the autonomy of robotic surgical systems and potentially reducing the cognitive load on surgeons, thus increasing surgical safety.
Building on these advances, the development of Efficient Class-Promptable Surgical Instrument Segmentation represents a further evolution in the field. Models like the SAM prompting system can be directed toward segment-specific classes of surgical instruments through simple prompts or minimal user input such as drawing a bounding box or a point prompt of the target object [86]. This capability is crucial in surgical settings, where the types of tools and the visual environment can vary significantly between procedures. Such models allow for rapid adaptation to new instruments without the need for extensive retraining, improving their practical utility in the operating room. This adaptability not only increases segmentation accuracy but also reduces computational demands, making real-time implementation more feasible [87].
Further enriching the field, advances in foundational models, particularly those based on Transformer architectures, have significantly enhanced surgical image segmentation. These models utilize self-attention mechanisms to effectively manage complex spatial relationships in medical imaging, complementing traditional techniques like CNNs [88, 31]. Innovations such as the SAM demonstrate the adaptability required for dynamic surgical environments, crucial for applications where rapid changes demand reliable model performance [43, 89]. Moreover, the integration of these advanced models into augmented reality systems provides surgeons with context-sensitive real-time information that improves both precision and safety during surgical procedures [90]. These would act as clinical decision support tools that highlight features that link real-time images with preoperative CT or MRI scans. For example, pre-operative CT detection of an enlarged lymph node with potential cancer cells in the mesentery next to the ileo-colic or middle colic artery could be highlighted in real time when performing an extended right hemicolectomy (removal of the right and transverse colon). This would help the surgeon achieve a clear surgical margin and reduce the risk of cancer recurrence.
The recent development of the Video-Instrument Synergistic Network (VISN) underscores a significant advancement in leveraging machine learning for robotic surgeries, as explained by Wang et al. [91]. VISN is specifically designed to handle the complexities of video-based instrument segmentation in this field. Taking advantage of synergistic interactions between video frames and instrument-specific features, it significantly enhances segmentation accuracy. VISN leverages the dynamic relationships and spatial-temporal consistency of surgical instruments within video sequences, achieving precise segmentation critical for automating robotic surgery. Its core strength lies in its ability to parse and integrate contextual information from continuous video feeds, seamlessly adapting to the variable and visually complex environment of robotic surgeries. This adaptation ensures high accuracy in instrument detection and segmentation despite challenges like variable lighting, occlusions, and the presence of blood and tissue fluids. VISN not only enhances the safety and efficacy of robotic-assisted surgeries by integrating into surgical workflows but also reduces surgeons’ cognitive load, potentially improving patient outcomes and surgical efficiency.
TABLE IV: State-of-the-Art Models for Surgical Instrument Segmentation of surgical tools and instruments within various surgical environments, as seen in robotic and laparoscopic surgeries.
Reference Dataset Model Technique
Garcia-Peraza-Herrera et al. [84] EndoVis15, NST, FFT FCN-8s+ Supervised
Attia et al. [92] EndoVis15 CNN + RNN Supervised
Garcia-Peraza-Herrera et al. [28] DVR ToolNet Supervised
Shvets et al. [93] EndoVis-17 Ternaus11, Ternaus16, LinkNet34 Supervised
Milletari et al. [94] EndoVis15 ResNet +, Conv LSTM Supervised
Ross et al. [95] EndoVis17 ResNet, U-Net Semi-Supervised
Pakhomov et al. [96] EndoVis15 ResNet+atrous Supervised
Islam et al. [97] EndoVis17 CNN+Residual Auxilary, Adversarial
Lee et al. [98] Private DCNN Weakly-Supervised
Fuentes-Hurtado et al. [99] EndoVis15 DeepLabv3+ Weakly-Supervised
Mohammed et al. [100] EndoVis17 StreoScenNet Supervised
Jin et al. [101] EndoVis17 MF-TAPNet Supervised, Self-supervised
Yu et al. [102] EndoVis17 Modified U-Net Supervised
Gonzalez et al. [103] EndoVis17,18 ISINet Supervised
Ni et al. [104] Cata7, EndoVis17 BarNet Supervised
Sahu et al. [105] Sim, Sim Cholec80, EndoVis15 DNN+, Ternaus11 Unsupervised Domain Adaptation
Zhao et al. [106] Davis16, EndoVis17,18, HKPWH AOMA Meta-Learning
Peng et al. [107] UW Sinus, EndoVis17 DeepLabv3+, MobileNet Active-Learning
Su et al. [108] Sinus Surgery CycleGAN Adversarial
Zhang et al. [109] Private, EndoVis17 Synthetic U-Net+PatchGAN Adversarial
Colleoni et al. [110] Custom FCNN Supervised
Lin et al. [111] Sinus Surgery LC-GAN Adversarial
TABLE V: Comprehensive Overview of State-of-the-Art Models for Instrument Detection for detecting surgical tools in dynamic surgical settings, showcasing their application on different datasets.
Reference Dataset Model Technique
Twinanda et al. [87] Cholec80, EndoVis15 EndoNeT Supervised
Teevno et al. [112] m2cai16 Teacher-Student Semi-Supervised
Zhang et al. [109] AJU-Set, m2cai16-tool- locations FasterRCNN+, Region proposal Network Supervised
Namazi et al. [113] M2CAI16, Cholec-80 RCNN Supervised
Hu et al. [114] m2cai16-tool AGNet Supervised
Mishra et al. [115] m2cai16-tool CNN+LSTM Supervised
Kurmann et al. [116] RMIT, EndoVis15 CNN Supervised
Sarikaya et al. [117] ATLASDione CNN + RPN Supervised
Vardazaryan et al. [118] Cholec80 FCN Weakly-supervised
Yoon et al. [119] Private Faster, CascadeRCNN Semi-supervised
Kondo et al. [120] Cholec80 CNN + Transformer Supervised
Alshirbaji et al. [121] Cholec80 CNN+LSTM Supervised
III-B Instrument Detection
Instrument detection in surgical settings is crucial for identifying surgical tools’ presence and location within images or video frames, setting the foundation for tracking movements and recognizing surgical activities. Most of the detection models leverage sophisticated frameworks like Faster R-CNN or YOLO for efficient multi-object detection in the dynamic environments of surgical scenes.
Twinanda et al. [87] introduced the EndoNet architecture, combining tool detection with phase recognition, exemplifying how multitask learning enhances surgical scene understanding by simultaneously tackling related tasks, thus improving the robustness and accuracy of tool identification and contextual understanding. In open surgery, DL models detect and localize tools within dynamic environments, which helps postoperative analysis and real-time decision making [122]. This capability not only improves surgical safety and efficiency by ensuring the correct use and placement of instruments, but also enhances surgical training by providing real-time feedback on tool handling.
Furthermore, integrating instrument detection with phase recognition, as further developed by Twinanda et al. [71], demonstrates the potential of combining multiple analytical tasks for an understanding of surgical workflow. Such systems contribute to automated documentation, improved situational awareness, and improved coordination of surgical procedures. Table V summarizes various state-of-the-art models for instrument detection, including techniques and models like EndoNet, Faster R-CNN, and Cascade R-CNN in datasets such as Cholec80, EndoVis15, RMIT, and m2cai16-tool. This table illustrates the evolution of instrument detection technology, emphasizing advancements towards sophisticated, multitask learning frameworks that promise greater precision in real-time surgical tool detection.
In addition to these applications, instrument detection technologies are instrumental in advancing robotic-assisted surgeries. Accurate detection and localization of surgical tools enable precise control and manipulation by robotic systems, thereby enhancing the surgeon’s capabilities and reducing the likelihood of errors. The ability to monitor tool positions in real-time also facilitates the implementation of safety protocols, such as preventing unintended movements and ensuring that instruments are used within their designated areas.
III-C Instrument Tracking
Instrument tracking in surgical environments involves continuous monitoring of surgical tools across video frames, critical for the effective operation of robotic systems and assistive technologies. This capability ensures accurate real-time localization of instruments, adapting seamlessly to surgeon actions and enhancing procedural safety and efficiency.
Table VI provides an overview of advanced state-of-the-art models for instrument tracking, including LinkNet and ST-MTL, which excel in complex tasks combining detection, segmentation, and tracking, essential for precise real-time instrument tracking in modern surgeries. Other models like U-Net and FCN are noted for their robust detection and accurate instrument positioning, which are crucial for maintaining performance in dynamic surgical settings. The table details the effectiveness of these models across diverse surgical environments, listing the specific datasets validated along with the tracking techniques used. This comprehensive information highlights how each model contributes to instrument tracking, offering insights into their operational efficiencies and applications.
TABLE VI: State-of-the-Art Models for Instrument Tracking of surgical instruments across video frames, highlighting their effectiveness in dynamic surgical environments.
Reference Dataset Model Technique
Zhang et al. [123] m2cai16-tool LinkNet Detection and Segmentation
Islam et al. [124] EndoVis17 ST-MTL Tracking, Segmentation
Allan et al. [125] Self Decision tree, Optical flow Features
Zhao et al. [126] Private CNN Detection
Sarikaya et al. [117] ATLASDione CNN+RPN, Fast RCNN Detection
Lejeune et al. [127] BRATS, EndoVis15,Cochlea U-Net Features
Du et al. [128] RMIT, EndoVis15 FCN Detection - Regression
Nwoye et al. [129] Cholec80 FCN+ ConvLSTM Detection
Traditional Tracking Methods
Traditional tracking methods such as the Kalman Filter are widely used in dynamic systems tracking due to their ability to predict state variable changes based on velocity and acceleration measurements. This method excels in environments with linear motion profiles and Gaussian noise, which can be reliably estimated under these specific conditions [130]. It uses a predictive model that iteratively updates state estimates, making it suitable for applications such as surgical tool tracking, where motions are typically linear. However, surgical instruments often display non-linear movement patterns with sudden changes in direction and speed, challenging traditional tracking methods. The assumptions of linearity and Gaussian noise integral to Kalman Filters may not apply in surgical contexts, leading to potential inaccuracies and impairing the filter’s effectiveness in complex surgical maneuvers [131].
DL-based Methods
To address the limitations of traditional tracking methods, advanced DL-based methodologies using neural networks have been developed, utilizing temporal consistencies and contextual data to improve the accuracy of the tracking. DL architectures like RNNs and Long Short-Term Memory (LSTM) networks excel in capturing complex temporal dependencies and adapting to the unpredictable dynamics of surgical tool movements [132]. These models trained on extensive datasets of annotated surgical activities, generalize well across different surgical scenarios and instrument types. A key development by Allan et al. [133] uses DL for real-time surgical tool tracking, combining CNNs with LSTM layers to effectively model the spatial and temporal aspects of surgical instruments. This hybrid approach adeptly manages rapid movements and visual occlusions, which are common challenges in surgical settings, enhancing both the functionality of surgical assistive technologies and the precision of robotic instrument control, thus reducing operational errors and improving outcomes.
Furthermore, recent studies have explored Transformer-based models, which are renowned for handling long-range dependencies within sequential data [134]. These models, with their scalability and adaptability, are well-suited for the complex data involved in surgical instrument tracking. Transformers employ attention mechanisms to selectively focus on relevant data segments, significantly enhancing tracking accuracy across extended periods and diverse surgical environments.
Sensor-Based and Integrated Tracking Systems
Sensor-based tracking systems augment vision-centric methodologies by incorporating data from electromagnetic trackers or Inertial Measurement Units (IMUs), adding dimensions of information that improve the robustness of the tracking [135]. These multimodal approaches counteract the limitations of vision-based systems, such as occlusions and variable lighting, by providing additional reliable sources of positional data. Integration of instrument tracking with detection and segmentation technologies is also critical, as it facilitates consistent and accurate representations of the surgical environment and allows functionalities such as automated tool re-identification, workflow analysis, and Augmented Reality (AR) overlays for surgical navigation [136]. Despite advances, tracking in dynamic and cluttered surgical settings remains challenging due to factors such as tool occlusions, variable lighting, and the presence of biological tissues, along with the need for real-time processing, which imposes directions on computational efficiency. Therefore, developing sophisticated models that balance tracking accuracy with computational speed is essential to maintain robust performance without disrupting surgical workflows.
(a) DeSAM [137]
(b) AdaptiveSAM [138]
(c) CycleSAM [139]
(d) Surgical SAM2 [140]
Figure 7: Segmentation tasks in surgical video analytics addressed by SAM variants: (a) precise tool localization via detection and segmentation integration (DeSAM); (b) improved segmentation in complex scenarios with task-specific prompts and multi-head attention (AdaptiveSAM); (c) temporal consistency through reference-based segmentation (CycleSAM); and (d) efficient video processing with memory-based frame pruning (Surgical SAM2). These adaptations highlight SAM’s versatility in tackling the challenges of surgical environments. Images adapted from corresponding references.
III-D Instrument Pose Estimation
Pose estimation in surgical settings involves traditional computer vision and modern deep learning-based methods. Traditional techniques using feature detection and geometric modeling often employ regression-based approaches such as Support Vector Regression (SVR) and Random Forests to predict pose parameters directly from images [141]. However, these methods may struggle with the variability of the surgical environment, such as instrument occlusions and various orientations. Deep learning advances have greatly improved the accuracy and robustness of pose estimation. Fully Convolutional Networks (FCNs) are used for articulated pose estimation, detecting multiple key points and their spatial relationships [128]. Furthermore, RNNs and Long Short-Term Memory (LSTM) networks also incorporate temporal information from video sequences to enhance real-time pose tracking [142]. These architectures manage temporal dependencies and motion dynamics, ensuring consistent pose estimations across frames, even in the presence of motion blur and occlusions.
Accurate instrument pose estimation is critical to improving various aspects of surgical practice. In robotic-assisted surgeries, it provides precise control of robotic arms, allowing surgeons to execute complex maneuvers with a reduced risk of instrument clash [143]. Pose estimation also supports AR systems in operating rooms by overlaying essential information onto the surgeon’s field of view, thus improving situational awareness and navigation [144]. Furthermore, in surgical training and simulation, it enables the creation of realistic simulators that offer real-time feedback on instrument handling and techniques, improving training effectiveness [126]. Despite its benefits, pose estimation in surgical settings faces challenges such as tool occlusions, variable lighting, and complex backgrounds. Real-time performance without compromising accuracy remains essential for practical application. Future research aims to enhance the robustness and adaptability of pose estimation systems through the use of unsupervised and semi-supervised learning techniques, the integration of multimodal data, and the incorporation of domain-specific knowledge to meet the unique requirements of surgical environments [135].
III-E Anomaly Detection and Safety Measures
Anomaly detection and safety measures use ML to identify deviations from normal operations that can indicate hazardous situations, in order to enhance surgical safety through alerts or corrective actions. Kayan et al. [145] explored how the integration of anomaly detection systems in robotic surgery can preemptively address risks by recognizing unusual tool movements or unsafe interactions.
III-E1 Complication Detection
Complication Detection focuses on identifying potential complications during surgeries, such as unexpected bleeding or improper instrument handling, using machine learning to analyze real-time surgical data. Ruan et al. [146] highlight continuous monitoring of the surgical process, integrating data from surgical tools and physiological monitors to enhance the detection capabilities of complications.
III-E2 Error Prediction
Error Prediction forecasts potential errors in the handling or functioning of surgical instruments to enhance safety and efficiency. Miao et al. [147] demonstrate how DL models can predict errors during robotic surgeries in real time, effectively reducing error rates by alerting surgeons to potential instrument misuse or failure before they compromise the procedure.
III-F Segmentation using Foundation Models
Building upon the foundation laid by traditional ML and DL techniques in surgical applications, this section shifts focus to foundational models, specifically the Segment Anything Model (SAM). In surgical applications, SAM is being refined for precise tool segmentation. This involves fine-tuning the model with surgical-specific datasets and incorporating domain-specific knowledge to enhance precision without sacrificing generalization across different surgical setups [148]. High-resolution surgical images and detailed annotations are used in training to help the model learn the nuanced features of surgical instruments and tissues [93]. Techniques such as transfer learning and domain adaptation adjust the model’s parameters for surgical contexts, with few-shot learning enabling performance improvements with limited data [87]. These advancements contribute to more precise segmentation outputs, which is crucial for real-time instrument tracking, surgical navigation, and augmented reality applications in the operating room [149]. Below are some variants of the SAM architecture as shown in Fig 7:
III-F1 Surgical-DeSAM
The Surgical-DeSAM explores the decoupling of the SAM’s components to better suit robotic surgical environments, where precision and reliability are paramount. By modularizing the segmentation tasks, DeSAM enhances the robustness and accuracy of instrument segmentation, which is critical for automated or semi-automated robotic surgeries [137]. Decoupling allows each module to be fine-tuned individually, enabling more precise control over the segmentation process and facilitating the integration of domain-specific knowledge relevant to robotic surgery [150]. In such robotic surgeries, accurate instrument segmentation is essential for tasks such as motion tracking, collision avoidance, and providing visual feedback to the surgeon [111]. The complexity of surgical scenes, combined with the variability of instruments and tissues, presents significant challenges for segmentation models. By decoupling SAM’s components, Surgical-DeSAM addresses these challenges by allowing for specialized processing of different aspects of the segmentation task. For instance, separate modules can handle the unique visual features of robotic instruments or adapt to varying lighting conditions in the surgical environment [95]. This modular approach also facilitates easier updates and maintenance of the system, as individual components can be improved or replaced without affecting the entire model [151].
III-F2 AdaptiveSAM
AdaptiveSAM [138] represents an evolution in the application of the SAM for surgical segmentation, focusing on dynamically adjusting the segmentation parameters in response to the specific needs of each surgical video frame. This adaptability ensures high precision and relevance of the segmentation output, adapted to the specific characteristics of the surgical regions [138]. In surgical environments, conditions such as lighting, tissue appearance, and instrument presence can change rapidly, making static segmentation models less effective [71]. AdaptiveSAM addresses this challenge by incorporating mechanisms that fine-tune the model on the fly, enhancing its responsiveness to the dynamic surgical scene. One of the key features of AdaptiveSAM is its ability to efficiently utilize computational resources while maintaining high segmentation accuracy [152]. By employing techniques like transfer learning and incremental learning, the model can adapt to new data without retraining from scratch [153]. This is particularly important in surgical settings where real-time performance is crucial and computational resources may be limited. AdaptiveSAM’s efficient tuning process allows it to provide immediate feedback to surgeons, assisting in tasks such as instrument tracking, tissue identification, and navigation [3].
III-F3 CycleSAM
CycleSAM introduces a novel approach that leverages cycle-consistent feature matching to enable one-shot learning within the SAM, allowing the model to perform accurate segmentation from a single annotated example. This capability is particularly valuable in surgical settings where obtaining comprehensive training data is challenging due to the scarcity of annotated images and the high cost of expert labeling [139]. In traditional segmentation models, a large amount of annotated data is required to achieve high accuracy, which is often impractical in medical contexts. CycleSAM addresses this limitation by using cycle-consistent adversarial networks to align features between the source domain (annotated examples) and the target domain (unlabeled surgical images) [154]. This method ensures that the learned features are robust and transferable, enabling the model to generalize from a single example to a variety of surgical scenes. The loss of cycle consistency enforces a bidirectional mapping between the domains, preserving the structural integrity of surgical images while adapting to new contexts [155]. This approach effectively mitigates the domain shift problem commonly encountered in medical image analysis, where variations in imaging conditions can significantly impact model performance [156]. By incorporating cycle-consistent feature matching, CycleSAM enhances SAM’s ability to perform accurate segmentation without extensive retraining or additional data collection. This one-shot learning capability is crucial in surgical environments, where real-time decision-making is essential and the availability of annotated data is limited. CycleSAM’s efficiency enables quicker deployment of segmentation models, facilitating applications such as surgical navigation, instrument tracking, and intraoperative guidance [3].
III-F4 Surgical SAM 2
Advancing the capabilities of the SAM, recent iterations focus on the real-time segmentation of surgical videos by employing efficient frame pruning techniques. These innovations reduce the computational load by selectively processing frames that contain significant changes or relevant surgical actions, enabling SAM to operate effectively in real-time surgical scenarios [140]. This is crucial in dynamic surgical environments where rapid processing is essential for assisting surgeons without disrupting the workflow. By minimizing latency and ensuring timely feedback, efficient frame pruning enhances the practicality of SAM in the fast-paced setting of an operating room. Efficient frame pruning works by identifying and discarding redundant frames that do not contribute new information, thereby optimizing resource utilization without compromising segmentation accuracy [157]. This approach allows the model to focus computational efforts on critical moments within the surgical video, improving both speed and efficiency. Researchers are also exploring the integration of motion detection algorithms and temporal consistency models to further enhance the performance of SAM in real-time applications [87]. These advancements make it feasible to deploy sophisticated segmentation models in surgical settings, facilitating better surgical navigation and instrument tracking and potentially improving patient outcomes.
III-G Endoscopic and Laparoscopic Tool Detection Datasets
The MICCAI Endoscopic Vision (EndoVis) challenge initiated in 2015 have played a crucial role in propelling AI research in endoscopic and laparoscopic tool detection. These challenges offer datasets for diverse tasks like segmentation, tracking, and classification. The availability of high quality datasets is vital in driving advances in the field. Starting with rigid surgical instrument segmentation and tracking in 2015 [158], subsequent challenges expanded to include binary and multi-class instrument segmentation, and by 2018, comprehensive scene segmentation tasks involving both robotic and non-robotic tools as well as anatomical structures were introduced [159]. The 2019 Robust-MIS challenge elevated this with a dataset of 30 surgical procedures aimed at improving robustness and generalization in binary and instance segmentation of surgical tools [160]. Additional focuses in recent years have included depth estimation in the SCARED sub-challenge, as well as domain adaptation and workflow recognition in the 2020 and 2021 challenges, respectively. The M2CAI16 Challenge presents two datasets for surgical workflow and tool detection: the m2cai-tool dataset [87] annotating the presence of seven surgical tools that feature 15 cholecystectomy videos from the University Hospital of Strasbourg where 10 are used for training and 5 for testing purpose. This dataset was later expanded to m2cai-tool-locations [161], adding spatial annotations for 2,532 frames to aid tool localization. The Cholec80 dataset, developed by Twinanda et al. [87], includes 80 cholecystectomy videos by 13 surgeons, with 86,000 annotated images for tool presence and surgical phase recognition, including tool bounding boxes for 10 videos. It was extended by the ITEC Smoke Cholec80 Image dataset [162], which adds 100,000 frames annotated to differentiate smoke from non-smoke scenarios, aiding research on smoke removal in surgical environments. The Kvasir-Instrument dataset [163] comprises 590 endoscopic frames from gastroscopies and colonoscopies, providing binary segmentation masks and bounding boxes for various instruments at resolutions from 720×576 to 1280×1024. Additionally, the ART-Net dataset [164] includes 29 laparoscopic hysterectomy procedure videos recorded in 1920×1080 resolution, featuring binary segmentation and geometric data for non-robotic instruments, enhancing its application in tool presence detection. All these details of the datasets are summarized in Table III.
IV ML/DL Applications in Surgical Workflow Analysis
Surgical Workflow Analysis critically employs ML to enhance the understanding and efficiency of surgical procedures by automating the recognition of phases and steps within surgeries. This analysis is key to improving efficiency, safety, and quality in surgical settings. Recent advancements in DL have significantly strengthened the ability to perform nuanced analyses of complex surgical activities. A detailed review by Demir et al. [165] highlights the crucial role of DL techniques in recognizing surgical phases and steps, employing advanced architectures such as CNNs, RNNs, and transformers to process sequential video data for identifying critical surgical actions and transitions. These models are particularly effective in complex environments rich in visual data.
The efficacy of these models in surgical workflow analysis is rigorously evaluated using benchmarks such as the HeiChole benchmark, which provides a standardized dataset from cholecystectomy procedures to validate and compare algorithms [166]. Such comparative analyses are vital for identifying superior computational approaches and fostering ongoing improvements in the field. In addition, there is an increasing emphasis on the application of these advanced analytical methods in specific surgical specialties, including open orthopedic surgery. This involves analyzing intraoperative video data to understand workflows, crucial for training, planning, and providing real-time operational support. Adapting DL models to the unique visual and procedural nuances of specialties like orthopedic surgery presents significant challenges, given the specialized instruments and complex movements involved.
IV-A Phase Recognition
Surgical phase recognition is a key aspect of surgical process modeling, with the aim of automatically identifying different stages of a procedure from video data. This capability is crucial to improving surgical training, automating documentation, and supporting real-time decision making in operating rooms. Recent advances in deep learning have notably advanced the accuracy and real-time functionality of phase recognition systems. Machine learning models, especially those employing temporal classification networks like LSTM networks, effectively recognize and segment different surgical phases from video data, structuring the surgical workflow to facilitate relevant information and tool availability at each stage and improving surgical efficiency and safety assessments. Twinanda et al. [87] introduced EndoNet, a DL architecture that performs phase recognition in surgical scenarios, demonstrating its utility in automating and improving understanding of surgical procedures.
A systematic review by Garrow et al. [24] examined the transition from traditional ML methods to advanced DL models in surgical phase recognition, tracing the evolution of increasingly sophisticated algorithms that utilize extensive datasets and complex model architectures [167]. This progression has significantly enhanced accuracy and real-time applicability. Early methods involved simpler classifiers like Support Vector Machines (SVM) and Hidden Markov Models (HMM), but these were often constrained by their linear decision boundaries and basic assumptions about data sequence dependencies [167]. The early advancements techniques favor DL methods, particularly CNNs, and RNNs, which effectively manage the spatial and temporal complexities of surgical videos [87, 168]. Moreover, the development of multi-stage temporal convolutional networks (TCNs), such as the “Tecno” model, offers a robust solution that bypasses the limitations of recurrent architectures by using temporal convolutions to capture sequential dependencies more efficiently [169, 170]. These TCNs are particularly adept at processing sequential data, making them ideal for real-time applications that require instant feedback and recognition during surgical procedures [169]. The comprehensive and fine understanding of these models significantly improves the automation of surgical documentation and aids in training surgical residents by offering objective and consistent phase identification [170].
IV-B Tissue and Organ Segmentation
Tissue and Organ Segmentation is vital in medical imaging, enabling precise identification and delineation of anatomical structures for presurgical planning and intraoperative guidance. DL models, especially U-Nets and Fully Convolutional Networks (FCNs), have dramatically improved the segmentation of medical images, offering high accuracy and real-time processing [29, 171]. The V-Net architecture, a three-dimensional variant of the U-Net developed by Milletari et al. [74], excels in volumetric segmentation, processing entire data volumes simultaneously to effectively capture spatial hierarchies of tissues. The use of volumetric convolutions and skip connections in this model helps to maintain segment boundary accuracy during up-sampling, which is crucial for surgeries requiring high precision. Furthermore, recent enhancements incorporate Generative Adversarial Networks (GANs) and attention mechanisms, which improve the model’s ability to differentiate between similar tissues and enhance boundary contrast in complex scenarios [172].
IV-C Anatomical Structure Recognition
Anatomical Structure Recognition is an essential component of advanced surgical systems, focusing on the accurate identification and classification of anatomical structures within the surgical field. This capability is crucial to improving the functionality of automated surgical systems and providing surgeons with real-time decision support. Using advanced ML techniques, particularly CNNs, these systems can efficiently recognize and label complex anatomical features, helping to improve the precision and safety of surgical interventions. Oh et al. [173] demonstrated the effectiveness of DL models in real-time recognition of anatomical structures, significantly improving both the accuracy and safety of surgical procedures. These models are trained on large datasets of annotated images, allowing them to develop a nuanced understanding of various anatomical nuances that are critical during operations.
IV-D Action and Task Recognition
Action and Task Recognition in surgery identifies specific actions and tasks within phases, crucial for training, performance evaluation, and protocol adherence. Advanced ML techniques, such as CNN and sequence modeling, analyze surgical videos to detect subtle surgeon activities, thus enhancing the monitoring and standardization of practices [174]. Some of the recent developments include a computer vision platform that uses ML algorithms to recognize actions and highlight crucial events during surgeries such as laparoscopic cholecystectomy, contributing to procedural standardization and training improvement [175]. Additionally, gesture recognition technologies in robotic surgery improve interactions between surgeons and robotic systems, enhancing efficiency and outcomes by providing intuitive control of surgical robots [176].
Autonomous Instruments Control uses ML to automate the control of surgical instruments, improve precision, and minimize human errors by replicating expert maneuvers, offering consistency in intricate tasks, and reducing surgeon fatigue [177, 178]. Anomaly Detection and Safety Monitoring employs ML to identify deviations from standard procedures, enhancing safety with real-time alerts and interventions, which are essential in high-stakes environments [166, 179]. It is imperative that these alerts do not overload clinical staff, resulting in “notification fatigue”. Striking a balance is key.
Augmented Reality (AR) and Navigation Assistance technologies use ML to overlay critical information directly on the surgical field and provide navigational signals during procedures, significantly impacting surgical planning and execution [180, 181, 3]. Error detection and feedback mechanisms use ML to monitor surgeries and provide corrective feedback in real time, thereby improving quality, safety and supporting surgical training by providing real-time guidance and post-procedure analysis to beginner surgeons [178, 87, 182].
IV-E Downstream Tasks using Multimodality
BERT is a foundational language model renowned for its proficiency in understanding and generating human language. In the surgical domain, BERT has been leveraged to improve various aspects of clinical practice, documentation, and decision-making processes. One of the primary applications of BERT in surgery involves tasks of natural language processing (NLP), such as the automatic extraction and interpretation of information from surgical reports, electronic health records (EHR) and operative notes [183]. By accurately parsing complex medical terminology and contextual nuances, BERT-based models facilitate the creation of structured data sets from unstructured textual data, which can be instrumental for clinical research and patient management.
In addition, BERT has been integrated into decision support systems to help surgeons make informed decisions during procedures. For example, by analyzing historical surgical data and outcomes, BERT models can provide predictive insights and recommendations tailored to specific patient cases, thus improving surgical planning and risk assessment [184]. Furthermore, BERT improves communication within surgical teams by enabling more efficient information retrieval and summarization, ensuring that all team members have access to relevant and up-to-date patient information in real-time. This approach to key information enables high performance within the team setting.
In the realm of surgical training and education, BERT-powered applications offer personalized learning experiences for surgical trainees. These applications can analyze trainee performance reports, identify areas for improvement, and suggest targeted educational resources, thus fostering continuous professional development [185]. Furthermore, BERT facilitates the development of intelligent virtual assistants that can support surgeons by answering questions, providing procedural guidelines, and managing administrative tasks, ultimately contributing to increased operational efficiency and reduced cognitive load during surgeries.
LLaMA is renowned for its natural language processing prowess and is enhancing surgical and healthcare operations through a series of specialized systems. LLaVA-Surg, as detailed by [186], uses LLaMA to integrate and interpret multimodal data from surgical procedures. This system helps surgical teams by providing dynamic support during operations, helping to identify crucial procedural steps and potential deviations, thus improving surgical safety and efficiency. Following this, Surgical-LLaVA focuses on deepening the understanding of surgical scenarios. Jin et al. [187] highlight how this system utilizes large language and vision models to process complex surgical data, thus improving the precision and responsiveness of surgical interventions.
Furthermore, LlamaCare by [188] leverages LLaMA’s capabilities to improve healthcare knowledge sharing. This system improves the dissemination of medical information and best practices throughout the healthcare community, facilitating better communication between healthcare providers and improving patient care through access to updated and comprehensive medical knowledge.
ChatGPT has also shown significant promise as an intraoperative and educational tool in various surgical disciplines, as highlighted in recent studies. Atkinson et al. [189] and Araji et al. [190] demonstrate how ChatGPT helps decision making during complex surgeries such as Deep Inferior Epigastric Perforator (DIEP) flap procedures and enhances learning experiences for surgical clerkships. These studies illustrate ChatGPT’s ability to deliver real-time, accurate responses and educational support, improving surgical safety and training outcomes. Extending its applications, the review by Goglia et al. [191] further emphasizes ChatGPT’s role in abdominal and pelvic surgery, showing its effectiveness in preoperative planning, intraoperative decisions, and patient communication. Together, these studies reflect a transformative shift towards integrating AI in surgical practices, underscoring ChatGPT’s potential to optimize surgical workflows, precision, and educational protocols while also highlighting the need for ongoing refinement and cautious integration into clinical settings.
IV-F Workflow Analysis Datasets
Workflow analysis in surgical settings benefits from a variety of specialized datasets, each contributing unique insights and challenges to the field. The ATLASDione dataset, compiled by Sarikaya et al. [117], includes 99 videos from robotic tasks performed by clinicians, annotated with tool bounding boxes, action types, durations, and surgeon’s skill levels. The dVPN data set from Ye et al. [192] provides 34,320 pairs of stereo images for training and 14,382 pairs for testing, improving stereo vision and depth estimation in robotic surgeries. The AutoLaparo dataset [193] contains 21 high-definition videos of laparoscopic hysterectomy procedures, annotated for tool and anatomy segmentation and workflow recognition. The NeuroSurgicalTools data set [158] offers 14 neurosurgery videos with detailed instrument annotations, addressing visual challenges like occlusions and reflections. The FetalFlexTool dataset by García-Peraza-Herrera et al. [84] provides ex-vivo fetal surgery images and videos, annotated for tool segmentation under varied conditions.
Additional synthetic and simulated data sets include the UCL data sets [194, 110], offering synthetic images and videos rendered to simulate various surgical scenarios. The Laparoscopic Image-to-Image Translation dataset [195] includes 100,000 images derived from CT scans, adapted to various visual styles for model testing. The Sinus Surgery Dataset [196] features videos of sinus surgeries annotated for tool segmentation in complex environments. Cata7 [197] focuses on cataract surgery, providing high-resolution images with detailed annotations, while the CaDTD data set [198] extends the CATARACTS data set [199] with semi-supervised learning techniques for cataract surgeries. Lastly, the RoboTool dataset [200] includes real and synthetic images for training in robust tool segmentation. These diverse data sets collectively improve the development and testing of machine learning models in various surgical specialties and conditions.
V ML/DL Applications in Surgical Training and Simulation
ML and DL are significantly enhancing surgical training and simulation, offering profound insights and enhancements in the understanding of surgical scenes. These technologies help create realistic simulations of surgical procedures, allowing trainees to practice and learn in a safe environment. ML and DL tools generate detailed visuals and provide information about patient anatomy and how surgical tools interact with tissues. This helps surgeons understand and navigate the surgical environment better, making it easier for them to perform successful operations. Overall, using ML and DL in surgical training helps improve a surgeon’s skills and knowledge, which is crucial for effective and safe surgery. Following are some of the key components that correspond to designing efficient tools and techniques.
V-A Key Frame Extraction
The extraction of key frames from surgical videos is a critical process for efficient video analysis, allowing the identification of important moments without the need to review entire video sequences. An advanced method, proposed by Ma et al. [201], uses a diverse and weighted dictionary selection algorithm to identify key frames based on their representativeness and uniqueness. This approach ensures that the selected frames capture critical phases of the surgery, enhancing their educational and analytical value. Another technique, described by Tan et al. [202], leverages large-scale DL models to extract sequential key frames, accurately summarizing surgical procedures. This method emphasizes the extraction of key stages of the surgery, providing a concise and informative visual summary, particularly useful for surgical training and procedural reviews.
V-B Tissue Classification
Tissue Classification is crucial in surgical procedures for tasks such as cancer excision, involving the differentiation of tissue types based on characteristics such as appearance, texture and morphology. Machine learning, particularly DL-based techniques, excels in analyzing surgical images and videos to classify tissues with high accuracy, aiding surgeons in making decisions during complex interventions. Cekic et al. [203] introduced a DL approach using CNNs to classify tissue types in surgical videos in real time, effectively distinguishing between healthy and pathological tissues. This precision is vital in oncological surgeries to ensure complete removal of malignant cells while conserving healthy tissue. The field has seen further advancements by integrating more sophisticated architectures like Residual Networks (ResNets) and Inception networks. These networks provide deeper and more complex models that capture a broader range of features from surgical images, significantly enhancing the accuracy of classification [204].
V-C Depth Estimation and 3D Reconstruction
Depth Estimation and 3D Reconstruction technologies play a crucial role in providing three-dimensional perspectives of the surgical field from two-dimensional images or video feeds, particularly in MIS. These technologies, leveraging ML models such as stereo vision algorithms and structured light approaches, enhance depth perception and spatial understanding, thus improving the precision and safety of surgical interventions. Guni et al. [205] highlighted the effectiveness of ML techniques in producing accurate 3D reconstructions that significantly aid in planning and executing complex surgical procedures. These models are trained on large datasets and identify and interpret depth signals to create detailed 3D maps of the operative area. Recent advancements include using DL frameworks like CNNs and Generative Adversarial Networks (GANs) to improve the accuracy and detail of 3D reconstructions. CNNs process large amounts of image data to detect crucial edges and contours for depth estimation, while GANs enhance detail in partially obscured areas [206]. Furthermore, the use of time-of-flight (ToF) cameras and laser triangulation with ML algorithms facilitates real-time 3D reconstruction during surgeries, providing immediate feedback for on-the-fly adjustments, crucial for successful surgical outcomes [207].
V-D Surgical Video Generation
The generation of realistic surgical video through ML models opens new possibilities for training, simulation, and research. A recent work designed SurGen [208], an innovative approach to surgical video generation which is a text-guided diffusion model designed to create detailed and accurate surgical videos from textual descriptions. This model leverages the capabilities of diffusion models, which construct images by gradually refining patterns of noise into structured visuals. By integrating text input, SurGen allows users to specify what the video should depict, making it a powerful tool to create customized training materials or to simulate surgical scenarios to study possible outcomes and strategies, as shown in Fig. 8.
(a) Gallbladder dissection.
(b) Clipping and cutting.
(c) Preparation for removal.
(d) Calot’s triangle dissection.
Figure 8: The use of SurGen [208], a text-guided diffusion model for surgical video synthesis, to generate key stages of gallbladder surgery. These images highlight the potential of foundation models to improve surgical scene understanding by generating realistic visuals for training and planning. Images adapted from [208].
Recent advancements in the field have shown that video generation models can also leverage temporal coherence and multi-modal data to enhance the realism of generated surgical procedures. For example, the work by Wang et al. [209] introduced a method for surgical video synthesis that maintains temporal consistency across frames, which is crucial for realistic simulations. Similarly, Yamada et al. [210] explored the integration of multi-modal inputs such as surgical instrument tracking data, further improving the fidelity of generated videos by aligning visual and kinematic information.
V-E Anomaly Detection
Anomaly detection in surgical settings is critical for ensuring patient safety and enhancing surgical precision. This technology leverages advanced ML techniques to identify deviations from normal procedures, which can indicate potential risks or complications that might occur when the position of the surgical dissection plane alters.
In recent developments, unsupervised learning techniques have been applied to robotic surgery through the use of deep residual autoencoders. These models learn to reconstruct normal surgical activities by training on large datasets of standard operations. During actual procedures, significant deviations from the expected reconstruction are detected and flagged as anomalies, highlighting potential issues. The unsupervised nature of this approach makes it particularly valuable in environments where labeled data is scarce or unavailable, making it an ideal tool for real-time monitoring of complex robotic surgeries [211]. This method ensures that deviations from expected surgical patterns are identified, improving surgical accuracy and patient safety.
This method focuses on capturing temporal and spatial abnormalities in endoscopic videos of the esophagus, as shown in Fig. 9. By employing CNNs that track changes over time, the system can detect subtle anomalies that may be indicative of esophageal diseases, such as early-stage cancers or dysplasia [212]. This method demonstrates how ML can improve diagnostic precision in challenging clinical contexts.
Figure 9: Detection results from complex endoscopic frames, highlighting challenges in identifying esophageal abnormalities in surgical scenes: (a) tool occlusion, (b) bubbles, (c) motion blur, and (d) fog. Red boxes indicate AI model predictions, and blue boxes represent ground truth annotations. Images from [212].
Moreover, [213] provides a comprehensive review of DL applications in the detection of medical anomalies, highlighting key methodologies, successes, and ongoing challenges. It underscores the importance of DL in enhancing the diagnostic capabilities of medical imaging technologies. The survey highlights several critical advancements in the field, such as the ability of DL models to automatically learn and extract intricate features from large medical datasets, significantly improving the accuracy and efficiency of anomaly detection compared to traditional methods. For example, CNNs have been widely adopted for detecting anomalies in medical imaging, particularly for identifying abnormalities in radiological scans, such as tumors in CT or MRI images. RNNs and LSTM networks, on the other hand, are particularly effective in sequential medical data, such as electrocardiograms (ECG) and electroencephalograms (EEG), where temporal dependencies are crucial to detect irregularities.
V-F Surgical Video Summarization
Surgical video summarization is crucial in the medical field for education, documentation, and preoperative planning, enabling efficient review of lengthy and complex surgeries. ML and DL techniques play a vital role in this process by identifying key events, actions, and phases in surgical videos to create concise summaries without losing essential information. One prominent method is the automatic summarization of endoscopic surgical videos, where systems like those developed by King et al. [214] automatically generate summaries by identifying critical surgical phases, condensing lengthy recordings into essential, manageable segments. This allows surgeons and trainees to quickly review procedures, capturing major events such as the start and end of critical phases. Zisimopoulos et al. [215] introduced another method using hierarchical DL models that capture both temporal and spatial information to accurately identify important segments. This technique uses a combination of convolutional and recurrent neural networks to provide contextually rich summaries.
Esteva et al. [216] suggest combining video summarization with real-time feedback systems to enhance surgery decision-making, potentially integrating with surgical robots or augmented reality platforms for real-time procedural guidance. Advanced techniques further refine surgical video summarization such as King et al. [217], they utilized object detection and Hidden Markov Models (HMMs) to segment and summarize skull base surgeries by highlighting key instruments and actions. In addition, live tags from surgical teams mark important moments, enabling collaborative summarization. A deep multi-scale pyramidal features network dynamically summarizes complex surgeries by capturing hierarchical structures in the content, suitable for procedures requiring multiple detail levels [218]. Furthermore, techniques such as deep feature matching and motion analysis specifically tailor summarization for wireless capsule endoscopy by focusing on areas of pathological interest [219]. These innovations collectively improve the field of surgical video summarization, providing powerful tools for education, documentation, and clinical review.
V-G Surgical Skills Assessment
Surgical skills assessment is a fundamental aspect of maintaining high standards in surgical training and practice, and recent advancements in ML have led to the development of innovative and objective methods for evaluating these skills. One such approach is the Contrastive Regression Transformer model, which is designed to assess surgical skills during robotic surgeries by analyzing video data. This model captures subtle movements and decision-making processes to provide a quantitative assessment that aids in improving surgical performance [220]. Additionally, video-based analysis of recognized surgical gestures and skill levels has been explored in the work by Wang et al. [221], where ML techniques are used to correlate specific actions with skill levels, offering a more detailed and interpretable evaluation of surgical proficiency. This method enhances the feedback given to surgeons during training, ensuring a more tailored and effective learning process. Furthermore, the link between a surgeon’s technical skills and patient outcomes has been well-established, as highlighted by Stulberg et al. [222], emphasizing the critical impact that high surgical proficiency has on improving clinical outcomes. These developments demonstrate the growing role of ML in refining surgical training and ensuring optimal patient care.
TABLE VII: Overview of Foundation Models in Medical Imaging across Various Modalities and Tasks. This table summarizes key methodologies, their model specifications, modalities, downstream tasks, and primary findings. The abbreviations here are Seg: Segmentation, Cls: Classification, Det: Detection, IE: Image Enhancement
Year Method Code Model Modality Downstream Task Key Finding
2023 MA-SAM [46] ViT (SAM) CT, MRI, Endoscopy Seg Improved robustness across modalities
2023 Endo-FM [223] ViT Endoscopy Seg, Cls, Det Surpasses SOTA with significant margins in pre-training and transfer learning
2023 VisionFM [224] - Multimodal images Cls Delivers expert-level diagnostic accuracy in ophthalmology, surpassing traditional models and specialists with its modality-agnostic approach
2023 Polyp SAM++ [225] ViT (SAM) Endoscopy Seg Advances segmentation using text-guided SAM, enhancing localization and accuracy, effectively handling complex colonoscopy images
2024 SP-SAM [226] ViT Endoscopy Seg Integrates part-level prompts with image embeddings, enhancing detailed structure segmentation, with minimal parameters
2024 Polyp-SAM[227] ViT (SAM) Endoscopy Seg Utilizes fine-tuned SAM for polyp segmentation, achieving state-of-the-art results with superior generalization, Dice scores
2024 Swinsam[228] ViT (SAM) Endoscopy Seg Integrates Swin Transformer decoder with SAM encoder, improving segmentation detail significantly, enhancing performance metrics
2024 SurgicalSAM[86] ViT (SAM) Endoscopy Seg Introduces efficient tuning for SAM with a class prompt encoder, achieving SOTA performance with reduced parameter complexity and improved pipeline simplicity
2024 USFM[229] - US Seg, Cls, IE Demonstrates adaptability across organs and tasks with minimal annotations, enhancing robust feature extraction with spatial-frequency dual masking
2024 LVM-Med[230] ResNet, ViT Multimodal images Seg, Cls, Det Sets new benchmarks in medical imaging by leveraging large-scale self-supervised learning and second-order graph matching to enhance feature learning and adaptability
2024 VoCo[231] Swin CT Seg, Cls Utilizes volume contrastive learning to encode high-level semantics, significantly enhancing performance on segmentation and classification tasks
2024 RudolfV[232] ViT(DINO-v2) Pathology Cls Utilizes pathologist-informed design and semi-automated data curation from a diverse dataset, significantly enhancing diagnostic accuracy and robustness across various benchmarks
2024 AFTerSAM[233] ViT (SAM) CT Seg Enhances SAM with Axial Fusion Transformer, improving 3D medical image segmentation by integrating intra-slice and inter-slice contextual information with minimal training data
2024 PUNETR[234] - CT Seg Introduces prompt tuning for efficient semantic segmentation, achieving significant parameter efficiency with robust performance on medical imaging datasets
V-H Vision Language Foundation Model
DINO [55] has shown promise in revolutionizing surgical imaging by enabling improved interpretation and segmentation of medical images through its self-supervised learning approach. In surgical settings, DINO’s ability to learn detailed visual representations from unlabeled surgical imagery can significantly improve the accuracy and efficiency of surgical planning and intraoperative guidance. For example, DINO has been applied to automated segmentation of tumor boundaries in real-time during oncological surgeries, helping surgeons achieve more precise excision and margin control. This technology also facilitates the development of advanced diagnostic tools that can automatically differentiate between healthy and pathological tissues, enhancing the surgeon’s ability to make informed decisions during complex procedures [56]. In addition, by improving the quality and usability of intraoperative images, such as in laparoscopic video feeds, DINO contributes to safer and more efficient surgical workflows. This capability is crucial for MIS procedures, where the clarity and detail of visual information directly impacts precision surgery and therefore reduced postoperative complications.
Recently, “Surgical-DINO,” an adaptation of the DINO model, has been specifically designed to enhance depth estimation in endoscopic surgery, demonstrating the adaptability of foundational models to specialized surgical tasks [57]. Moreover, the SURGIVID project leverages DINO’s self-supervised learning framework for annotation-efficient surgical video object discovery, further highlighting its utility in reducing the labor-intensive process of video annotation in surgical training and analysis [56]. Furthermore, DINO’s capabilities could be instrumental in analyzing datasets like EgoSurgery-Tool, which focuses on detecting surgical tools and hands from egocentric perspectives in open surgery videos, potentially enhancing the training and performance of AI models in recognizing and interacting with complex surgical environments [235]. The adoption of DINO in these surgical applications not only streamlines the surgical process but also opens possibilities for more adaptive and responsive surgical systems, potentially reducing the incidence of complications and improving overall surgical outcomes[236].
CLIP and other vision-language foundation models (VLMs) are also reshaping the landscape of medical imaging by effectively linking textual descriptions to visual data. This capability is pivotal for diagnosing radiological images, enhancing surgical planning, and facilitating educational initiatives. The zero-shot learning capability of CLIP introduced by Radford et al. [63] allows it to recognize and categorize medical conditions across various imaging modalities, such as X-rays and MRIs, without direct training on specific medical datasets. This cross-modal understanding is especially beneficial in scenarios where large annotated datasets are scarce. Building on this application, Kerdegari et al. [237] demonstrated how foundational models could be adapted to improve the detection and categorization of pathologies in specialized medical imagery, such as endoscopy images for the diagnosis of gastric inflammation (gastritis). This study illustrates CLIP’s potential to enhance diagnostic accuracy in complex imaging scenarios.
Further expanding the scope of foundational models in healthcare, Sun et al. [238] provides a comprehensive examination of medical multimodal foundation models in clinical diagnosis and treatment. Their research focuses on various applications, challenges, and future directions of these technologies in healthcare. They highlight how these models integrate diverse data types by combining visual, textual, and possibly even genomic information to offer more nuanced diagnostics and treatment strategies. This integration facilitates a deeper understanding of patient data, crucial for personalized medicine and advanced treatment planning. Zhao et al. [239] also contribute to this body of knowledge by surveying the deployment of CLIP in medical fields, revealing its transformative impact from routine diagnostics to complex surgical interventions. Table VII provides a detailed overview of these foundational models, outlining their specifications, modalities, and key advancements in enhancing precision and robustness across various medical imaging and clinical tasks.
V-I Visual Question Answering
TABLE VIII: Summary of recent advancements in Visual Question Answering for surgical applications
Paper Key Findings Dataset Used Model Type Application Area
Surgical VQA [240] Introduced transformer-based Surgical-VQA for integrating visual-text information in training. Surg-VQA Transformer-based Surgical training
LV-GPT [241] Developed LV-GPT for enhanced word-vision token sequencing in VQA. Surg-QA Language-Vision GPT Robotic surgery assistance
SSG-VQA-Net [242] Used scene graph-based model for scene-aware reasoning in surgical VQA. SSG-QA Scene Graph Network Scene understanding
Surgical-VQLA++ [13] Improved VQA robustness via adversarial contrastive learning. Surg-LA Contrastive Learning Error detection
Surgical-LVLM [243] Applied specialized perception blocks for complex surgical scenes. Surg-LVLM Vision-Language Model Complex surgical environments
Surgical-VQLA [244] Used gated vision-language embedding for localized answering. Robo-Surg Gated V-L Embedding Robotic surgery
Surgical-LLaVA [187] Tailored large language and vision model for multi-modal chat in surgery. LLaVA-Surg LLaVA Model Multimodal assistance
LLaVA-Surg [186] Advanced multimodal surgical assistant capabilities with Surg-QA. Surg-QA Vision-Language Video question answering
GP-VLS [245] General-purpose VQA model with comprehensive surgical datasets. GP-VLS General Purpose V-L Model Surgical evaluation
CAT-ViL [246] Co-attention gated model for effective training in robotic surgery. Surg-CAT Co-Attention Model Surgical training
Visual Question Answering (VQA) in surgery uses AI to analyze visual data from surgical videos, enhancing decision-making and educational experiences for medical professionals. This application interprets dynamic visual content in real time, deepening understanding of complex surgical procedures, and serves as a crucial tool for training and operational assistance. The Surgment system pioneers this field with advanced segmentation techniques, enabling specific queries about visual elements in surgical videos, such as identifying tools or recognizing procedural actions, thereby enriching video-based surgical training [247]. The Surgical-VQA method utilizes transformer architectures, adept at processing video streams, to provide precise, contextually relevant answers, enhancing real-time surgical decision-making [240].
Further extending these capabilities, the Language-Vision GPT (LV-GPT) model integrates visual data processing with traditional GPT-2, using a vision tokenizer and vision token embeddings designed for surgical VQA. This model outperforms existing frameworks on datasets like the Endoscopic Vision Challenge Robotic Scene Segmentation 2018 and CholecTriplet2021, setting new benchmarks for AI-driven VQA applications in surgery [241]. Furthermore, VQA technologies significantly aid postoperative analysis, allowing surgical teams to review complex procedures with AI-enhanced visual data analysis, potentially improving surgical techniques and outcomes. Integrating VQA with augmented reality (AR) and virtual reality (VR) technologies could further revolutionize surgical training and operative workflows by providing AI-generated annotations and insights during procedures, enhancing surgeons’ field of view, reducing errors, and improving outcomes. Recent advances in VQA for surgical applications include transformer-based models and sophisticated systems that improve the robustness and precision of VQA tasks in complex scenarios. These developments underscore the increasing importance of VQA systems in surgical training, decision support, and procedural guidance, promising to transform surgical education and operations by improving the cognitive capabilities of surgical systems.
Table VIII summarizes recent advancements in VQA for surgical applications, showcasing a variety of models, datasets, and key findings that improve tasks such as scene understanding, error detection, and surgical training. It highlights innovative approaches like transformer-based models and graph networks, which enhance the interpretability and accuracy of surgical VQA systems.
V-J Surgical Video Retrieval
Efficient retrieval of surgical video content is vital in education, training, and clinical review, allowing users to quickly access relevant video segments based on specific queries. One notable method involves unsupervised feature disentanglement, which separates and identifies key features within surgical videos to improve the accuracy of retrieval systems in MIS [248]. This unsupervised approach improves the ability to search large volumes of video data (analogous to a comprehensive surgical video library) without relying on extensive labeling, thus streamlined access to critical information during surgical review and training [249].
TABLE IX: Comparative analysis of deep learning models for real-time segmentation and instrument identification in various surgical procedures.
Reference Procedure Dataset Model Application Modality Task/Tool Results
[250] Cholecystectomy m2cai16-tool + Cholec80 3D DenseNet+GCN Instrument Classification Images Tool 90.2% + 90.13% (mAP)
[251] Cholecystectomy Cholec80 LSTM Phases Boundary Detection Images Task 48 s (MAE)
[252] In-vitro experiments JIGSAWS + NPA VGG16 Gesture Segmentation Images + Kinematics data + events Task 86.3% + 89.4% (acc)
[253] Cholecystectomy Cholec80 ResNet18 Fine-grained activities Images Task 24.8% (acc)
[254] In-vitro experiments JIGSAWS Dense CNN Trajectory Segmentation Images + Kinematics data Task 70.6% (mAP)
[255] Cholecystectomy + Colorectal surgery Cholec80 VGG-50 + LSTM Phases Recognition Images Task 89.2% (acc)
[162] Cholecystectomy Cholec120 ResNet-152 Surgery Time Prediction Images Task 460 s (MAE)
[256] Gynecology NPA GoogleNet Phases Recognition Images Task 79.6% (AP)
[257] Nine different surgeries NPA VGG16 + LSTM Surgery Type Recognition Images Task 75% (acc)
[258] Colorectal Surgery EndoVis 2015 3D FCNN Instrument Detection Images Tool 85.1% (Dice)
[259] Porcine Procedures EndoVis 2017 ResNet18 Instrument Segmentation (Binary + Parts) Images Tool 89.6% + 76.4% (Dice)
[260] In-vitro experiments Atlas Dione + EndoVis 15 Stacked Hourglass Instrument Detection Images Tool 98.5% + 100% (mAP)
[261] Gynechologic surgeries NPA GoogleNet Anatomical Structures Classification Images Task 78.1% (acc)
[262] Sleeve Gastrectomy EndoVis 2015 Xception Surgical Scenario Segmentation Images Task 98.44% (Dice)
[263] Colonoscopy EndoVis 2015 VGG16 Polyp Detection Images Task 80% (Dice)
[98] Phantom % Porcine Tissue NPA LinkNet-152 Binary Instrument Segmentation Images Tool 88.9% (Dice)
[264] Laparoscopic Surgery Surgical videos with RGB images Multi-task CNN with U-Net encoder Blood accumulation detection and tool semantic segmentation Video Tool segmentation and event detection 81.89% + 90.63% (Dice)
[265] Minimally Invasive Surgeries EndoVis2017, EndoVis2018 Custom network with classification module Instance segmentation of surgical instruments Image Instrument segmentation Improves SOTA by at least 12 points on EndoVis2017
[86] Surgical instrument segmentation EndoVis2018, EndoVis2017 SurgicalSAM with prototype-based class prompt encoder Segmentation using SAM with enhanced class prompting Image Instrument segmentation SOTA performance with minimal tunable parameters
[91] Robotic Surgery Custom surgery video datasets VIS-Net with Graph-based Relation-aware Module Referring video instrument segmentation using text Video Instrument segmentation based on language descriptions Significantly outperforms existing methods
[266] Laparoscopic Surgery Dresden Surgical Anatomy Dataset with 13,195 images DeepLabv3, SegFormer Anatomy segmentation Image Semantic segmentation of anatomical structures 0.23 to 0.85 (IoU); outperforms human experts
[267] Laparoscopic Surgery DSAD SegFormer, DeepLabv3 Anatomy segmentation, enhance real-world applicability Image Semantic segmentation of anatomical structures Improved IoU, accuracy, precision, recall, F1 score, specificity, Hausdorff Distance, and ASSD
[268] Pituitary Surgery Custom Dataset PitSurgRT with HRNet Real-time landmark detection and semantic segmentation of critical anatomical structures Video Instrument segmentation, anatomy localization Improved landmark detection mean error; IoU increased by 4.39%
[269] Prostate Surgery Intraoperative Videos of RALP Reinforcement U-Net Real-time semantic segmentation in robotic prostatectomy Video Segmentation of instruments, bladder, prostate, seminal vesicle–vas deferens Dice score 0.96, 0.74, 0.85, 0.84 respectively; IoU 0.77
[270] Laparoscopic and Partial Nephrectomy CholecSeg8k, Private Dataset Spatio-temporal network Enhance temporal consistency in video semantic segmentation Video Surgical scene segmentation Improved temporal consistency and mean IoU by 1.30% and 4.27% on datasets
[271] General Surgery EndoVis2018, EndoVis2017 ASI-Seg with SAM Audio-driven instrument segmentation with surgeon intention understanding Image Semantic segmentation and intention-oriented segmentation 82.37% + 71.64% (IoU)
V-K Concluding Insights and Implications
This review has delved into a broad range of methodologies and applications in the realm of surgical video analysis and instrument detection, as illustrated in Table IX. The exploration underscores the rapid progress in deploying FMs, such as advanced neural networks, to enhance surgical outcomes and support real-time decision making.
These technologies have evolved from basic instrument classification to sophisticated dynamic phase recognition, showcasing the substantial potential to increase the precision and efficiency of surgical procedures. Adopting state-of-the-art neural architectures and segmentation techniques has particularly enhanced the sensitivity and specificity of these models, making them essential tools in the operating room. This progression demonstrates how foundational models not only refining current surgical practices but also setting the stage for future innovations in medical technology.
V-L Computational Complexity of FMs in Surgical Applications
FMs have revolutionized the field of surgical AI by offering robust generalization capabilities and exceptional performance across a variety of tasks [272]. However, these models typically require vast amounts of data and considerable computational resources for training, making it challenging for deployment in clinical environments. This section explores effective strategies to mitigate the computational demands of foundational models, ensuring their practical applicability in surgical settings.
Foundational models like SAM [43] and CLIP [63] encompass diverse tasks and modalities, which require extensive training on large datasets. This extensive training enables them to perform well across different tasks but also makes them resource-intensive. In surgeries, where responses need to be immediate and accurate, the heavy computational demand of these models can be a significant limitation. For this purpose, several innovative techniques have been developed to adapt foundational models to be more resource-efficient, making them suitable for real-time surgical applications:
  • Adapters: Adapters are small modules added to a pre-trained model that can be adjusted to new tasks. They require changing only a few parts of the model, which reduces the computational load significantly. In surgery, adapters can help tailor models to recognize specific instruments or actions using limited data without needing extensive retraining.
  • Low-Rank Adaptation (LoRA): LoRA modifies a model’s deep structures in a way that reduces the amount of data they handle at once, which cuts down on the computing power required. This technique is useful for refining models to perform specific tasks like VQA and grounding for intricate surgical contexts [243].
  • Prompt Tuning: This method tweaks the inputs given to the model to guide it toward a particular function using the model’s existing knowledge base. This approach is computationally light and can be used to adjust models for specific tasks, such as analyzing surgical videos, without extensive reprogramming [273].
  • Knowledge Distillation: This process involves training a smaller, more manageable model to mimic a larger one. The smaller model retains much of the larger model’s effectiveness but uses less computational power, making it better suited for use in surgeries where fast processing is crucial [274].
  • Quantization and Pruning: These techniques reduce the model’s size and speed up its operations. Quantization decreases the precision of the numbers the model uses, and pruning removes parts of the model that have little impact on its performance [275]. Both adjustments help the model run faster and more efficiently, which is essential in a surgical setting.
VI Open Issues and Future Research Directions in Surgical Scene Understanding
Surgical scene understanding has made significant strides, yet numerous challenges and open issues persist that require ongoing research and innovative solutions, as shown in Fig. 10. Addressing these issues is critical for advancing the field and realizing the full potential of AI-driven surgical tools. This section categorizes these challenges into specific areas that are crucial for the development and deployment of surgical AI systems.
Figure 10: An illustration of potential challenges and future research directions surgical AI development.
VI-A Technical Challenges
VI-A1 Data and Annotation Constraints
One of the most persistent challenges in surgical video analytics is the scarcity and variability of annotated surgical data. Developing robust semi-supervised [276], weakly supervised [277], and unsupervised learning [278] algorithms that can effectively leverage unlabeled data remains an essential area of future research [279]. Such methods reduce the dependency on large annotated datasets [280], which are costly and time-consuming to produce. Furthermore, fostering collaborations across medical institutions can enable secure sharing of surgical videos and annotations under stringent privacy regulations. These efforts will expand the availability of diverse datasets for training and validating AI models, improving their robustness and generalization.
VI-A2 Integration with Clinical Workflows
Although AI models hold great promise for surgical scene understanding, their integration into clinical workflows remains limited due to technical and ergonomic barriers. Future research should prioritize the development of user-friendly interfaces and real-time analysis tools that seamlessly integrate with existing surgical equipment and protocols [281]. Consideration must also be given to minimizing the ergonomic and cognitive load on surgeons to ensure that AI systems serve as effective tools that enhance, rather than disrupt, surgical performance.
VI-A3 Robustness and Generalization
AI models used in surgical scene understanding often struggle with robustness and generalization due to the variability in surgical environments, practices, and equipment across hospitals and geographic regions [282]. Adaptive models that can generalize to new settings without extensive retraining are critical for broader adoption. Promising techniques such as domain adaptation and federated learning offer potential solutions to address these challenges and enhance the reliability of AI models in diverse surgical contexts.
VI-A4 Explainability and Trust
The “black-box” nature of many deep learning models used in surgery poses significant challenges in terms of explainability and trust [283]. Future research should focus on incorporating explainable AI principles that provide transparent and interpretable insights into AI decision-making processes [284]. The enhancement of the explainability of the model will be instrumental in gaining the trust of medical professionals and meeting regulatory requirements, ensuring that AI systems are reliable and safe for use in clinical settings.
VI-A5 Advancing Robotic Surgery
AI-driven insights derived from surgical videos have the potential to revolutionize robotic surgery by enabling more intuitive and responsive control systems [285]. Future research should focus on a deeper integration of AI analytics with robotic systems to support more sophisticated instrument handling, decision making, and task execution. This will improve the precision and safety of robotic-assisted surgeries while enabling novel applications in complex procedures.
VI-B Ethical and Regulatory Considerations
VI-B1 Balancing Automation and Human Skill Development
The increasing integration of AI and robotics in surgery, while enhancing precision and efficiency, raises critical concerns about the atrophy of human skills. As Beane notes in the article “Today’s Robotic Surgery Turns Surgical Trainees into Spectators” [286], surgical trainees are often relegated to passive observers in robotic surgeries, significantly limiting their opportunities to develop essential cognitive and motor skills. This overreliancen automation risks creating a generation of surgeons proficient in operating machines but ill-prepared for handling emergencies or unanticipated complications. To mitigate this, AI systems should prioritize human augmentation by enabling active trainee participation and providing real-time feedback, fostering skill development alongside technological progress [287] [288].
VI-B2 Ethical Considerations
As AI becomes increasingly integral to surgical workflows, addressing ethical challenges such as patient safety, data privacy, and bias mitigation will be crucial [289]. Future research should emphasize the development of robust ethical guidelines and frameworks to ensure that AI systems align with principles of fairness, accountability, and transparency. These guidelines will play a critical role in fostering trust and adoption among clinicians and patients.
From an ethical point of view, the shift towards automation introduces questions of accountability, transparency, and sustainability. As highlighted in The Skill Code [288], surgical AI must be designed to preserve human expertise rather than replace it, ensuring that technology complements the role of the surgeon. Over-standardization and dependency on AI could undermine creativity and adaptability in surgery, diminishing the human element essential for patient-centered care. Addressing these challenges requires a human-centered approach to AI that balances innovation with ethical responsibility, safeguarding both the art and science of surgery.
VI-B3 Regulatory Considerations
The rapid advancement of surgical AI requires robust and evolving regulatory frameworks to ensure safety, transparency, and compliance. Existing guidelines, such as AI as a medical device legislation from regulatory bodies such as the FDA and EU MDR, emphasize risk management, performance validation, and post-market surveillance [290]. To align with these standards, surgical AI systems must ensure data security, model interpretability, and clinical efficacy while incorporating mechanisms for continuous learning and adaptation.
Collaboration between AI researchers, medical professionals, and regulatory bodies is essential to establish comprehensive standards that balance innovation with rigorous oversight. These regulations should prioritize patient safety, fairness, and accountability to mitigate biases and unintended consequences. By fostering trust and ethical responsibility, such frameworks will support the sustainable integration of AI into surgical workflows, ensuring both efficacy and safety in clinical applications.
VI-C Multi-modal Data Fusion and Integration
VI-C1 Multi-modal Data Fusion
Combining information from surgical videos with other types of data, such as patient medical records, real-time sensor data, and histopathological images, can provide a more comprehensive understanding of surgical scenes [291]. Multimodal AI systems capable of leveraging this holistic integration will not only improve surgical precision and outcomes, but also enable more personalized and context-aware decision making.
VI-C2 Personalized Feedback
AI systems have significant potential to enrich surgical training and simulation through personalized feedback mechanisms. Future advancements could focus on the development of virtual reality (VR) and augmented reality (AR) platforms powered by AI to simulate diverse surgical scenarios and complications. These platforms can provide objective assessments and customized feedback to enhance surgeon training skills [292].
VII Conclusion
The deployment of machine learning and deep learning technologies in the realm of surgical scene understanding marks a significant paradigm shift in minimally invasive surgery. By leveraging sophisticated models like CNNs, ViTs, and FMs, these technologies will enhance the precision of surgical interventions, improve real-time decision-making, and ensure patient safety. However, foundational models have evolved the way these models operate as they achieve more diverse tasks and employ different applications like accurate segmentation and recognition of surgical instruments and tissues, which in turn reduces surgical risks and improves outcomes. Challenges remain in the field, such as handling diverse and complex surgical scenarios, ensuring model transparency, and integrating these technologies into clinical workflows. Future research should focus on developing more adaptable, efficient, and ethically aligned models that can be seamlessly integrated into clinical settings. Advances in the field will continue to push the boundaries of what is possible in surgery, potentially leading to groundbreaking surgical techniques and transforming surgical training and patient care. Moreover, the excitement about AI’s impact on healthcare is growing at the highest levels, with global healthcare leaders increasingly recognizing its transformative potential.